### v2

# Reasoning Abilities of Large Language Models: In-Depth Analysis on the Abstraction and Reasoning Corpus

Abstract.
The existing methods for evaluating the inference abilities of Large Language Models (LLMs) have been results-centric, making it difficult to assess the inference process. We introduce a new approach using the Abstraction and Reasoning Corpus (ARC) dataset to evaluate the inference and contextual understanding abilities of large language models in a process-centric manner. ARC demands rigorous logical structures for problem-solving, making it a benchmark that facilitates the comparison of model inference abilities with humans. Experimental results confirm that while large language models possess weak inference abilities, they still lag in terms of logical coherence, compositionality, and productivity. Our experiments highlight the reasoning capabilities of LLMs, proposing development paths for achieving human-level reasoning.
Contents
- 1 Introduction
- 2 Preliminaries
- 3 Evaluating the Inferential Capabilities of LLMs Using the ARC Benchmark
- 4 Discussion
- 5 Conclusions
- A Supplementary Analysis
- B Experimental Detail
1. Introduction
Recent Large Language Models (LLMs) have demonstrated performance levels close to that of humans, but experimental results showed that they lacked planning ability through thought or reasoning (Bubeck et al., 2023). Consequently, a key question in recent language model research is: Can LLMs think? To address this question, new benchmarks for measuring reasoning abilities such as MathVista (Lu et al., 2024), Bongard-Logo (Nie et al., 2020), and Raven (Zhang et al., 2019) have been proposed. Among them, the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) emerged to be one of the representative benchmarks for assessing reasoning abilities. As shown in Fig. 1 below, each task in ARC consists of 2–5 demonstration example pairs and a test example input grid. The goal is to infer rules from given example pairs and apply them to the test example. Input and output grid size can vary from a minimum of to a maximum of , with each grid having up to 10 different colors.
The ARC remains an unsolved challenge despite its seemingly simple content and evaluation methods. It demands a high level of abstraction and multiple reasoning steps, reasons why conventional deep learning techniques have not achieved success. The best-performing models to date have only achieved an accuracy of 3-40% (Lab42, 2024), while LLMs (GPT-4, PaLM) have shown an accuracy of around 10-20% (Mirchandani et al., 2023). Compared to the average human accuracy of 80% (Johnson et al., 2021), these results suggest significant differences in reasoning and abstraction capabilities between humans and LLMs. However, in-depth research into how LLMs reason and how their reasoning differs from humans is lacking. This has led to calls for a shift from a results-focused evaluation to a more nuanced analysis of the process (Chang et al., 2024; Huang and Chang, 2023; Xue et al., 2023; Allen-Zhu and Li, 2023), indicating a need for a new perspective that evaluates reasoning abilities based on the process rather than just the outcome.
To overcome the limitations of result-oriented analysis in artificial intelligence, this study adopts an existing theory on what constitutes a human’s reasoning ability. According to the Language of Thought Hypothesis (LoTH) (Fodor, 1975), human reasoning encompasses three essential characteristics: Logical Coherence, the ability to maintain consistency in reasoning; compositionality, the capability to construct complex ideas from simpler components; and productivity, the capacity to formulate an indefinite number of thoughts or solutions using a finite set of elements.
While attempts to evaluate logical coherence, compositionality, and productivity have existed before (Bubeck et al., 2023; Valmeekam et al., 2024), there were limitations in that the definitions of each component varied across papers and existing benchmarks showed insufficient performance in assessing each aspect. This study differs from previous research in two key ways: 1) by redefining concepts borrowed from psychology to fit the field of computer science, and 2) by evaluating all elements through the visual reasoning benchmark ARC. To achieve these goals, we have designed three separate experiments:
Logical Coherence: LoTH identifies two types of coherence. These are inferential coherence — the ability to apply logical reasoning across related instances coherently — and semantic coherence — the ability to maintain logical coherence in the reasoning process and results (Fodor and Pylyshyn, 1988). To verify both types of logical coherence, we augmented each solved ARC task with 100 similar test examples and evaluated the LLM’s performance on these related instances. Additionally, we analyzed the solution processes, identifying cases where correct answers were derived from flawed reasoning, to measure the LLM’s semantic coherence.
Compositionality: Compositionality refers to a system’s capacity to express one proposition is inherently linked to its ability to express related propositions (Fodor and Pylyshyn, 1988). In this study, we define compositionality as the ability to combine given semantics. Therefore, to evaluate compositionality, it is necessary to verify whether semantics can be combined as desired. Consequently, this study provided LLMs with step-by-step functions and examined whether they could identify the appropriate functions to solve ARC problems. Subsequently, we conducted an additional analysis to determine if the LLM could accurately predict the results from the given step-by-step functions and to understand the reasons for the failure.
Productivity: Productivity refers to the ability to infinitely create unseen expressions by combining a limited set of semantics (Fodor and Pylyshyn, 1988). However, it is difficult to quantitatively measure whether one can make an infinite number of unseen expressions. Therefore, previous studies have evaluated productivity by assessing whether rule-compliant unseen expressions can be created (van der Velde et al., 2004; Lake and Baroni, 2018; Hupkes et al., 2020). Similarly, in this study, to evaluate the ability to generate unseen expressions, we examined whether unseen ARC tasks that comply with the rules could be generated when given a set of functions.
As a result, we have confirmed that the current level of LLM possesses a basic understanding of images and is capable of simple types of compositional object manipulations. However, compared to human reasoning abilities, LLMs lag in three areas: 1) It is not inferentially and semantically coherent. 2) Its logical reasoning abilities, especially in a step-by-step manner, are weak. 3) It struggles with understanding and generating unseen representations under complex constraints.
Finally, this study summarizes and presents recent trends proposed to address the weaknesses in abstraction abilities and reasoning capabilities. Analyzing the reasoning abilities of LLMs according to the components of human reasoning and discussing how to enhance each component represents a differentiated approach from previous research. It offers a fresh perspective for measuring and advancing the reasoning capabilities of LLMs in the future.
2. Preliminaries
This section aims to explain why we chose the LoTH perspective and ARC before starting a detailed evaluation of LLM’s reasoning capabilities. First, we will look at existing definitions of reasoning abilities and show why LoTH is useful in the perspective of measuring intelligence in Section 2.1. Then, in Section 2.2, we show that the ARC is an appropriate benchmark for studying LLMs from the perspective of human reasoning, as it 1) utilizes abstract semantics that can be generalized, and 2) is easy to modify.
2.1. Limitation on Assessing Reasoning Ability of LLMs
Efforts to evaluate LLMs’ capabilities continue, underscoring strengths in image and text generation. Especially, analysis confirms LLMs possess elements of a World Model (Gurnee and Tegmark, 2023), indicating potential in inference tasks. Despite these capabilities, challenges in reasoning are noted (Valmeekam et al., 2024), with errors such as distortion and incomplete reasoning highlighted (Li et al., 2024). Research suggests these reasoning abilities can improve through methodological adjustments. Furthermore, studies indicate that complex compositionality remains challenging (Dziri et al., 2023).
The divergent claims about the reasoning abilities of LLMs stem from result-centric measurement methods. Turing was the first figure to shift the approach to inference towards consequential direction (Turing, 1950). Subsequently, Wiener (Wiener, 1950), McCulloch and Pitts (McCulloch and Pitts, 1943), and Rosenblatt (Rosenblatt, 1958) shifted to studying methods for measuring performance rather than focusing on the process. Recently, Chollet attempted to quantify inference abilities from a consequential perspective (Chollet, 2019). However, these studies all focus on what reasoning can achieve using a result-oriented approach, without specifying the elements that constitute reasoning ability. West et al. (West et al., 2024) raised concerns about evaluating the reasoning ability of LLMs from a consequentialist perspective, as the generation capability of LLMs may not necessarily depend on comprehension abilities.
Therefore, a new perspective is needed to evaluate AI’s inference processes; the Language of Thought Hypothesis (LoTH) enhances discussions by integrating reasoning components with quantitative metrics. LoTH posits that inference involves manipulating mental representations, a view that dominates the philosophy of mind due to its explanatory power over logical coherence, compositionality, and productivity observed in human cognition. These mental representations are believed to have a compositional syntax and combinatorial semantics. Our study compares with prior works and evaluates LLMs’ inference capabilities through the LoTH, marking progress by assessing aspects like logical coherence, compositionality, and productivity.
Logical coherence, compositionality, and productivity of LLMs have been evaluated separately in previous research. The ability of LLMs to perform deductive reasoning and maintain consistency in that reasoning has been considered an important evaluation metric. In this context, logical coherence is generally defined as the ability to construct coherent logic when solving specific problems (Zhang et al., 2024b). Meanwhile, whether LLMs possess compositional ability is one of the emerging questions. Given the lack of definition for compositionality, existing studies have not yet reached a unified opinion. Nevertheless, there have been attempts to define compositionality as the ability to understand and combine complex expressions, and to evaluate it by providing step-by-step prior knowledge necessary for solving problems requiring deductive reasoning (Lake and Baroni, 2018). Lastly, productivity is the ability to create new expressions from limited resources. With the emergence of generative AI, discussions are ongoing about what specific abilities or processes should be considered as productivity, and how to evaluate them. Some studies have judged this based on the accuracy and efficiency of the output generated using limited resources (van der Velde et al., 2004; Hupkes et al., 2020). However, these existing attempts have limitations in that the evaluation criteria and definitions vary across papers, and they have not compared the elements of reasoning to human reasoning processes.
Exploring deeper into the three perspectives of LoTH offers strong justification for improving reasoning capabilities. These principles help in developing the ability to process information and solve tasks similar to human reasoning. Logical coherence ensures LLMs can reason without contradictions, compositionality allows LLMs to adapt known knowledge to new scenarios, and productivity enhances LLMs’ capacity to generate results based on given rules. Thus, adopting these perspectives of LoTH aids LLMs in achieving more human-like reasoning, enabling them to address complex problems with innovative and valid results.
2.2. Advantages of using ARC as Reasoning Benchmark
In exploring benchmarks suitable for evaluating inference abilities through the lens of the Language of Thought Hypothesis (LoTH), the Abstraction and Reasoning Corpus (ARC) emerges as a compelling candidate. ARC has two strengths compared to other reasoning benchmarks. First, it aligns with the LoTH perspective in that it can be solved through combinations of semantics. Second, it allows for easy task modification and generation, enabling flexible changes to objectives.
2.2.1. Core Properties of ARC
An important characteristic of the ARC benchmark is its requirement for extracting compositional semantics and combining them to resolve problems. This feature distinguishes ARC from other benchmarks and necessitates sophisticated problem-solving approaches. Two key research findings support this assertion:
Importance of Semantic Information: Studies demonstrating enhanced performance through supplementary information underscore the significance of semantic content in ARC task resolution. In a notable experiment, the integration of graph-represented object information resulted in a substantial accuracy increase, nearly doubling the success rate (Xu et al., 2023b). This marked improvement emphasizes the critical role of semantic information in effectively addressing ARC challenges.
High Abstraction Level of ARC: Comparative analyses reveal ARC’s higher abstraction level relative to other benchmarks, as evidenced in Table 1. Chollet, the benchmark’s proposer, contends that conventional feature extraction methodologies are insufficient for ARC, given its demand for complex shape interpretation and transformation process comprehension (Chollet, 2019). These findings highlight the necessity of employing advanced strategies capable of interpreting the abstract content integral to ARC problem-solving.
These observations confirm the importance of developing and implementing approaches that can effectively extract and utilize the complex, abstract information essential to solving ARC challenges. The unique nature of ARC necessitates a shift from traditional problem-solving paradigms toward more sophisticated, semantically-aware methodologies. Future research in this domain should focus on creating algorithms that can not only process visual information but also infer underlying rules, patterns, and transformations.
| Input Shapes | Hidden Rules | Target Tasks | Specific Challenges | |||||
| Dataset | Geometric Shapes | Abstract Shapes | Explicit Rules | Abstract Rules | Classify | Generate | Domain Transfer | Extrapolate |
| ARC (Chollet, 2019) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Mini-ARC (Kim et al., 2022) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| 1D-ARC (Xu et al., 2023b) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| MC-LARC (Shin et al., 2023) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Sandia (Matzen et al., 2010) | ✓ | ✓ | ✓ | |||||
| Synthetic (Wang and Su, 2015) | ✓ | ✓ | ✓ | |||||
| G-set (Mańdziuk and Żychowski, 2019) | ✓ | ✓ | ✓ | |||||
| RAVEN (Zhang et al., 2019) | ✓ | ✓ | ✓ | ✓ | ||||
| PGM (Barrett et al., 2018) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Hill et al. (Hill et al., 2019) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| G1-set (Mańdziuk and Żychowski, 2019) | ✓ | ✓ | ✓ | ✓ | ||||
| S1-set (Mańdziuk and Żychowski, 2019) | ✓ | ✓ | ✓ | ✓ | ||||
| MNS (Zhang et al., 2020) | ✓ | ✓ | ✓ | |||||
| VAEC (Webb et al., 2020) | ✓ | ✓ | ✓ | ✓ | ||||
| DOPT (Webb et al., 2020) | ✓ | ✓ | ✓ | ✓ | ||||
2.2.2. Flexibility in benchmark adaptation
Despite its simple rules, ARC remains a challenging benchmark with relatively low accuracy. LLMs achieve 15% accuracy (Qiu et al., 2024), traditional program synthesis models reach 26% (Wind, 2020), while human average accuracy is 80% (Johnson et al., 2021). To address this challenge, various ARC variants have emerged, focusing on reducing either abstraction complexity or reasoning dimensions.
Three notable ARC variants are:
1D-ARC (Xu et al., 2023b): Reduces dimensionality from 2D to 1D, simplifying complexity while retaining core knowledge. It effectively addresses object cohesion challenges, resulting in high LLM accuracy (about 90%).
MC-LARC (Shin et al., 2023): Adopts a multiple-choice format, transitioning from generative to selection tasks. GPT-4 showed strong performance (approximately 75%).
Mini-ARC (Kim et al., 2022): Limits grid size to 5x5, simplifying input while maintaining 2D generative characteristics. Performance remains challenging, similar to original ARC (around 15%).
These variations demonstrate ARC’s transformation flexibility and support the necessity of composition in solving ARC tasks. As shown in Fig. 3, MC-LARC and 1D-ARC reduced reasoning step complexity, while Mini-ARC focused on reducing image complexity. The performance differences among these variants imply that reducing the need for complex transformation combinations can significantly improve results, highlighting the importance of combinatorial syntax in solving ARC.
In summary, the ARC emerges as a compelling benchmark for evaluating inference abilities through the lens of the LoTH. ARC’s core strength lies in its requirement for extracting and combining compositional semantics to solve tasks, aligning well with the LoTH perspective. This is evidenced by improved performance when additional semantic information is provided, such as object information represented as graphs. The various ARC variants (1D-ARC, MC-LARC, and Mini-ARC) demonstrate their flexibility and adaptability for different experimental purposes. The significant performance differences among these variants, particularly the high accuracy in 1D-ARC and MC-LARC compared to the original ARC and Mini-ARC, support the necessity of combinatorial syntax and sequential transformations in solving ARC tasks. Furthermore, ARC’s high level of abstraction and reasoning complexity, as seen in its low accuracy rates for both LLMs and traditional models compared to human performance, underscores its value as a challenging benchmark. These characteristics collectively validate the rationale for using ARC in this study as an effective tool for exploring inference abilities in the context of the LoTH.
3. Evaluating the Inferential Capabilities of LLMs Using the ARC Benchmark
To evaluate whether LLMs possess inferential capabilities, one could compare these capabilities to human reasoning. As explained in Section 2.1, according to the Language of Thought Hypothesis (LoTH), human reasoning can be broadly divided into three main components: Logical Coherence (Section 3.1), Compositionality (Section 3.2), and Productivity (Section 3.3). We utilized Abstraction and Reasoning Corpus (ARC) to examine each aspect of the reasoning capabilities of LLMs from the perspective of LoTH.
3.1. Capability of LLMs 1: Logical Coherence
3.1.1. Motivation
Section 3.1 aims to evaluate the logical coherence of Large Language Models (LLMs). This is a fundamental aspect of the Language of Thought Hypothesis (LoTH), which considers coherence in two senses: inferential coherence and semantic coherence (Fodor and Pylyshyn, 1988). Semantic coherence refers to the ability to maintain logical consistency in the process and results of reasoning. Inferential coherence, on the other hand, is a system’s ability to consistently apply a specific type of logical inference across all relevant instances, given it can perform that inference in some cases. These concepts are crucial in human cognitive processes and relevant to the rule inference required in ARC tasks.
Our initial experiments primarily focused on measuring semantic coherence by evaluating whether the results produced by LLMs logically followed their problem-solving steps. This was done using various prompt techniques such as Chain of Thought (CoT) (Wei et al., 2022), Least to Most (LtM) (Zhou et al., 2023), and Tree of Thought (ToT) (Yao et al., 2023), similar to previous ARC solving attempts (Xu et al., 2023b; Mirchandani et al., 2023). We compared the coherence levels these different prompting strategies achieved, aiming to identify which techniques yielded the most semantically coherent results across diverse problem-solving scenarios. However, recognizing the limitations of this approach in addressing inferential coherence, we introduced supplementary experiments using augmented ARC tasks. These tasks, created through the Re-ARC program (Hodel, 2024), allow us to assess how consistently LLMs can apply logical patterns across variations of originally solved problems, providing a more comprehensive evaluation of their logical reasoning capabilities.
3.1.2. Comparison Across Prompting Techniques
The perceived deficiency in LLMs’ logical reasoning has been a recurrent critique, with direct attempts at solving ARC tasks yielding success rates below 10% (Mirchandani et al., 2023). To mitigate this, enhancements in LLMs’ logical reasoning are being pursued through prompting techniques like Chain of Thought (CoT) (Wei et al., 2022), Least to Most (LtM) (Zhou et al., 2023), and Tree of Thought (ToT) (Yao et al., 2023). These strategies have been shown to leverage LLMs’ reasoning capabilities (Wang et al., 2019) effectively and have the advantage of allowing for a more transparent analysis for humans, as they involve a step-by-step reasoning process. Therefore, in this experiment, we assess the impact of these prompting strategies on LLMs’ logical coherence by solving ARC tasks.
We applied three major prompting techniques – Chain of Thought (CoT), Least to Most (LtM), and Tree of Thought (ToT) – to solve 100 ARC evaluation tasks using the GPT-4-32k model. Each technique was tested in five iterations. ARC tasks follow a few-shot learning paradigm, requiring the model to discern task rules from given example pairs and apply them to test examples. The CoT method enhances reasoning performance by utilizing a chain of thought, providing examples of ARC task solutions in the prompt. Similar contextual information was provided for LtM and ToT. LtM decomposes tasks into manageable steps and executes them sequentially, while ToT generates multiple candidates at each step post-decomposition, selecting the best candidate through voting before proceeding to the next step.
| Iteration | CoT | LtM | ToT |
| 1 | 11% (3%) | 6% (4%) | 7% (3%) |
| 2 | 10% (2%) | 7% (4%) | 4% (1%) |
| 3 | 10% (5%) | 6% (3%) | 7% (2%) |
| 4 | 10% (4%) | 4% (2%) | 7% (4%) |
| 5 | 10% (6%) | 5% (2%) | 6% (2%) |
| Average | 10.2% (4.0%) | 5.6% (3.0%) | 6.2% (2.4%) |
Comparing ARC accuracy across prompts, CoT demonstrated superior performance. Table 2 presents the results of applying LtM, CoT, and ToT to 100 randomly selected tasks from the ARC evaluation set. The experiment was repeated five times, with the percentage of correct answers included for each iteration. CoT achieved approximately 10% accuracy, while LtM and ToT showed about 6% accuracy. The superior performance of CoT may be attributed to the cumulative error propagation in ToT and LtM, where small mistakes in one step of their multi-step answer generation process can lead to compounded errors in subsequent steps. Based on these results, we exclusively used the CoT prompt in subsequent experiments.
However, when we checked the correctness of the solution process, all three prompting techniques showed low accuracy, with no significant difference at around 3%, as indicated in parentheses. These results demonstrate that while there may be differences in accuracy depending on the prompting technique, there is little variation in semantic coherence. It’s also important to note that both the results and processes fall far short of the average human accuracy of 80%. These findings suggest that LLMs lag behind humans in terms of logical coherence. To analyze the specific reasons for this, we conducted follow-up experiments. Section 3.1.3 analyzes inferential coherence, which is one aspect of logical coherence, while Section 3.1.4 examines the semantic coherence of LLMs through case studies.
3.1.3. Inferential Coherence of LLMs
In our second experiment, we designed a test to evaluate the inferential coherence of LLMs. Inferential coherence refers to the consistency in reasoning processes between the problem-solving steps and their outcomes. To assess this, we examined whether the LLM could solve problems similar to the ARC tasks it had previously solved successfully.
Fig. 5 provides an overview of the entire experiment. We began by using GPT-4o to solve test examples from 400 ARC tasks,111We used 400 tasks from the ARC training set. Unlike the previous experiment, we utilized the training set instead of the evaluation set because only the training set can be augmented through Re-ARC. repeating this process five times to identify which problems the model could consistently solve. For tasks that the LLM solved correctly at least once out of the five attempts, we employed Re-ARC (Hodel, 2024) to generate 100 additional test examples for each task. These augmented test examples are methodologically identical to the original tasks regarding their solution approach. We reasoned that if the LLM truly possesses inferential coherence and applies consistent logic in problem-solving, it should be able to solve all 100 of these augmented examples. This approach allowed us to rigorously test the LLM’s ability to generalize its problem-solving strategies across similar, but not identical, tasks.
Fig. 6 shows the results. Fig. 6(a) is a graph that represents all five iterations for the number of cumulatively correct augmented test examples. It’s important to note that the graphs for all five iterations show an exponential decrease. This graph trend doesn’t vary significantly between iterations, demonstrating low coherence regardless of the iteration. Fig. 6(b) shows the distribution of accuracy, averaged over five iterations, on 100 augmented test examples for each task. The notable points in this graph are that more than two-thirds of the tasks, 54 tasks, are concentrated below the average accuracy of 0.2, and the accuracy of the 8-th, which corresponds to the top 10%, is about 0.6. These figures demonstrate that LLMs have low inferential coherence for most ARC tasks.
3.1.4. Case Study: Semantic Coherence of LLMs
Finally, we analyzed how LLMs solved tasks in the two experiments described in Section 3.1.2 and Section 3.1.3. When evaluating not only the answers but also the process for the three prompts CoT, LtM, and ToT, we found that regardless of the prompt, the accuracy was about 3%, indicating that correct answers were being derived from incorrect processes, as shown in Fig. 7.
To solve the task, one needs to 1) identify objects within the input grid, 2) count the number of black squares inside each object, and 3) extract the object with the largest number of black squares. However, CoT, LtM, and ToT attempted to solve this task in incorrect ways. For CoT, objects appearing in the input grid were sorted first, then the object in the middle was selected. Even though CoT got the answer right, the method of sorting the objects was not comprehensible. For LtM and ToT techniques, it was understood that a specific object needed to be selected from the given input grid to solve the task. However, they mistakenly recognized objects from the test input grid. These solutions share a commonality in that they fail to explain a logically consistent rule between the provided examples of different training inputs and outputs. In other words, regardless of prompting techniques such as CoT, LtM, or ToT, LLMs have yet to demonstrate logical coherence in discovering a single rule that applies consistently across the examples provided to solve the task.
The inconsistency of inferring correct results from incorrect processes was also observed in the second experiment conducted on the training set. Upon analyzing the natural language explanations for the 83 tasks solved at least once out of 400 training tasks, we found that in 35 of these cases, the solutions proposed by the LLM could not produce the correct answer. This finding suggests that LLMs lack semantic coherence regardless of the prompting technique or the problem at hand. In other words, LLMs are deriving outcomes unrelated to their reasoning process, as evidenced by generating correct answers from incorrect solutions.
Nevertheless, in Section 3.1.3, we identified eight tasks that the LLM could solve with an accuracy of 0.6 or higher. As shown in Fig. 8, these eight tasks consists of simple solutions such as mirroring, color mapping, and partial grid copying. These tasks shared a common characteristic of conceptual simplicity, utilizing only one of the four prior knowledge domains included in ARC: objectness, goal-directedness, numbers and counting, and basic geometry (Chollet, 2019). For the 17 tasks that required the use of two or more prior knowledge domains, the LLM failed to solve any of the 100 augmented problems. The fact that the LLM could not solve any of the augmented problems, despite having solved the original ones, suggests that LLMs are not semantically coherent and may even indicate potential data leakage.
This comprehensive analysis demonstrates that while LLMs can solve certain simple pattern recognition tasks, they struggle with more complex reasoning that requires the integration of multiple concepts. The inability to coherently apply rules across augmented test examples, coupled with the generation of correct answers through incorrect reasoning, highlights significant limitations in both the inferential and semantic coherence of current LLM systems when tackling abstract reasoning tasks like those presented in ARC.
3.1.5. Conclusion
In Section 3.1, we evaluated the logical coherence of LLMs by solving 100 ARC tasks using three different prompting techniques. Our results, showing accuracies ranging from 4% to 12%, demonstrate variability in reasoning performance depending on the prompting approach used. Additionally, when experimenting with GPT-4o on 400 training tasks, the LLM showed a high accuracy of 20%.
However, through an in-depth qualitative review, we demonstrated that the LLM’s results may not be logically coherent. For the augmented test examples (100 for each solved task), the LLM only managed to solve 8 test examples showing performance above 60%. Furthermore, we found that for 35 out of the 83 solved tasks, nearly half, of the LLM provided incorrect solution processes that could not derive the correct results. This analysis suggests that the LLM has failed to achieve human-level logical coherence.
The results of this study align with previous research asserting that logical problem-solving remains challenging for LLMs alone. One study (Wang et al., 2023a) found that LLMs can generate logically consistent reasoning with CoT prompting, even when their reasoning steps are flawed. Another study (Zhang et al., 2024a) showed that LLMs struggle with accurate self-reflection in tasks like mathematical reasoning and translation. Additionally, research (Tyen et al., 2024) revealed that LLMs often fail to detect errors in intermediate steps, exposing flaws in their reasoning process. While these studies suggest that providing more context or enforcing stronger self-reflection might improve logical reasoning (Wei et al., 2022; Wang et al., 2023a; Zhang et al., 2024a), our findings indicate that these challenges persist, suggesting the issue may not be simply a lack of information about the problem.
3.2. Capability of LLMs 2: Compositionality
3.2.1. Motivation
In Section 3.2, we investigate compositionality, the second concept of LoTH.222While the Language of Thought Hypothesis principally uses the term ‘systematicity’, this study employs ‘compositionality’ as used in Fodor’s paper. This choice is made because compositionality encompasses a broader concept than systemicity. Compositionality refers to the ability to generate complex linguistic expressions given simpler ones (Fodor and Pylyshyn, 1988). This characteristic allows individuals to effectively tackle more complex tasks by breaking sub-tasks down into simpler steps, supporting the notion that humans can solve more complex tasks when faced with them. Strong compositionality enables the resolution of complex tasks and facilitates transparent descriptions of the process, which is also an important aspect from the perspective of LLMs.
This section uses ARC to test the compositionality of LLMs. There are previous studies that have tested a model’s compositionality by providing functions in the prompt that can be combined to solve tasks, and then checking if the model can solve them (Sinha et al., 2024). Similarly, in this study, we also provide step-by-step functions, which we refer to as DSL (Domain Specific Language), and then conduct experiments to verify whether they can solve ARC tasks. Additionally, to understand why tasks might not be solved, we conducted further experiments on the model’s comprehension of these functions. Therefore, we verify whether LLMs understand the meaning of the functions provided for ARC tasks and whether they can combine the functions appropriately to produce the desired results. The result of this experiment indicates that while LLMs sufficiently understand the functions and their relationship with images, their ability to decompose and combine functions to achieve the desired outcome is weak.
3.2.2. Compositionality of LLMs
In the first experiment, to measure compositionality, we provided LLM with information about DSL and asked them to solve given ARC tasks. Fig. 9 illustrates the structure of the entire experiment. If an LLM possesses sufficient compositionality, it should be able to select appropriate DSLs and their arguments for a given goal. However, in cases where the LLM failed to choose the correct DSL, we divided the conditions further to identify the cause. These conditions were whether the LLM understood the goal and whether it understood the solution process. To analyze the results according to each condition, four types of experiments were conducted: 1) given only DSL, 2) given correct output along with DSL, 3) given human descriptions to ARC test examples along with DSL, and 4) given both correct output grid and human descriptions along with DSL. Providing the correct output grid demonstrates compositionality based on knowing or not knowing the goal while providing human descriptions shows the impact of natural language descriptions on compositionality. All experiments were repeated 10 times each for 158 tasks.333The 158 tasks correspond to those that can be solved within 10 steps using the given DSL, out of the total 800 publicly available ARC tasks. Lastly, to establish a baseline, we also conducted human experiments. Seven participants were constrained to solve the tasks using only the same DSLs given to the LLMs.
Each DSL was given as a Python function. In this experiment, we used 19 types of DSL capable of solving ARC tasks. The prompts commonly included a brief explanation of ARC, DSL function code with comments, DSL usage examples, demonstration examples of tasks, inputs for the test examples, and object information of the test inputs. Object information is one of the crucial parameters in solving ARC tasks, which is why it was added to the prompt. We used the PnP algorithm (Park et al., 2023) to extract object information from ARC tasks. The LLM returned a JSON-formatted string representing the chosen DSL and arguments at each step, which was used to verify whether the LLM reached the correct test output with an appropriate combination of DSL and arguments. We used the most recent model, GPT-4o, for this experiment.
| w/o Human Description | w/ Human Description | Accuracy of Human | |
| w/o Test Output | 3% (0.93) | 8% (0.97) | 86% |
| w/ Test Output | 9% (0.96) | 14% (0.96) | X |
The experimental results are shown in Table 3. An average accuracy of 9% was observed when the test output was provided, and an average accuracy of 3% was observed without the test output. Similarly, it is an interesting finding that compositionality strengthens when human explanations are included in the prompt. As shown in Table 3, this improvement occurs at a similar rate to the test output. Cronbach’s alpha measurements for each experiment showed consistency in responses, with all four experiments scoring above 0.7. However, the series of results fell significantly short of the 86% accuracy corresponding to human performance. These low accuracy rates suggest issues in the LLM’s DSL composition process.
3.2.3. Analysis of compositional failures resulting from DSL misinterpretation
The issue is that the average accuracy described in Table 3 doesn’t solely reflect compositionality. When we use a DSL to solve ARC tasks, we can think about the likelihood of choosing the right DSL for each step in two parts: 1) How well LLMs understand the DSL: This is reflected in how accurately it can predict the next grid when given the DSL instructions. 2) How necessary each predicted grid is in creating the final solution: This relates to how well the steps fit together to solve the task. The overall chance of picking the correct DSL for all steps depends on both of these factors working together. To solve a task, all DSLs must be correct for 10 steps. Reflecting this, we can estimate as shown in Eq. 1 below. In this equation, represents the number of steps, represents the number of tasks at step , represents the single-step accuracy, and represents the difficulty of composition for each task. We assumed that the LLM’s compositionality could vary depending on the information provided to the LLM and the task.
To determine the task accuracy considering only the compositional difficulty, we must estimate the value when . Therefore, we conducted an additional experiment, as shown in Fig. 10 to verify the probability of not finding an appropriate DSL due to the inability to predict the output grid when selecting a DSL.
In the additional experiment, we checked how accurately the LLM could generate the correct output grid when given the DSL and ARC input grid. The experiment was conducted on 158 tasks, with each task repeated 10 times. The correct DSL and its argument chain for ARC tasks, created by humans while solving the tasks, were provided to the LLM. We prioritized using the solution with the shortest step length among the solutions provided by the human solvers. If the LLM can predict all subsequent output grids when given a specific DSL, it should be able to correctly produce the output grid regardless of the step length, since both the input grid and DSL were provided.
The LLM’s accuracy according to the number of DSL steps needed to solve the task is shown in Fig. 11. We can observe a tendency for accuracy to decrease as the number of steps required to solve the task increases. From the above results, we used a weighted average to calculate in Eq. 2. Eq. 2 shows the formula used to calculate p. In this equation, represents the single-step accuracy, represents the number of tasks at step , and represents the accuracy at step . Based on this, we estimated the single-step accuracy to be 81%.
Below Table 4 shows the expected modified accuracy when p corresponding to 0.8 is adjusted to 1. This table reflects only compositionality, confirming that nearly 30% of tasks can be solved when given the test output and human description. It should be noted that when both the correct answer and the natural language description leading to the answer are provided, there is a consistent increase of about 10% point. This suggests that each element positively influences the average compositional difficulty of the problem, represented as in Eq. 1.
| w/o Human Description | w/ Human Description | |
| w/o Test Output | 5% | 15% |
| w/ Test Output | 17% | 29% |
3.2.4. Case Study: Enhancement of Compositionality through Human Descriptions
One notable observation was the enhanced compositionality when human descriptions of problem-solving methods were included in prompts. To investigate how LLMs could solve tasks with human descriptions, we analyzed the solution processes of 13 additional tasks solved when human descriptions were provided. Results indicate that human descriptions facilitate task input abstraction and action abstraction, thereby improving problem-solving capabilities. For instance, as illustrated in Fig. 12, LLMs fail to recognize patterns in the correct output without descriptions; however, they immediately identify patterns such as an ‘X’ shape with descriptions. These findings suggest the potential to enhance LLMs’ reasoning performance by incorporating abstracted task information.
3.2.5. Conclusion
In Section 3.2, experiments using ARC and DSL were conducted to measure the compositionality of LLMs. The results led to three conclusions. First, LLMs were able to predict the output grid when DSL was applied to the input with an average accuracy of about 81%. However, as the length of steps increased, the prediction rate decreased, which appears to be due to cumulative errors. Second, when not given the correct answer, LLMs selected the correct DSL only 3% of the time, indicating a lack of ability both in inferring rules to predict the correct output grid and in selecting the appropriate DSL to reach the expected output. Finally, when human descriptions were added, the accuracy in choosing DSL increased to a level similar to when the correct answer was provided. Analysis of this process suggested that this improvement was due to linguistic abstraction of the ARC task and DSL combinations.
Previous studies have emphasized LLMs’ limitations in combining simple elements to create new meanings, revealing struggles with compositionality. One study shows Transformers exhibit significant performance drops when tested on new function combinations, indicating challenges in systematically generalizing knowledge (Hupkes et al., 2020). Another study introduced datasets like SADE to evaluate LLMs’ ability to process visual and textual information, suggesting they still struggle with tasks like understanding negations and grasping complex content (Ma et al., 2024). A further study examined how well LLMs can break down complex instructions or build them from simple ones. These found that while LLMs improve at understanding simple tasks by learning complex ones, they struggle with complex tasks when starting from simpler ones (Yang et al., 2024). These findings across studies point to ongoing challenges in LLMs’ ability to connect simple and complex elements, highlighting their compositionality limitations.
3.3. Capability of LLMs 3: Productivity
3.3.1. Motivation
In Section 3.3, we investigate the third concept of the Language of Thought Hypothesis (LoTH): productivity. Productivity refers to the ability to generate unseen representations based on observed data (Fodor and Pylyshyn, 1988). This characteristic enables humans to imagine diverse situations from a single phenomenon, facilitating efficient learning without the need for repetitive data exposure. Similarly, when endowed with this ability, LLMs are expected to excel in unseen tasks, making productivity a crucial function of essential reasoning. The capacity to generate new pairs within a constrained set of rules is particularly valuable for solving ARC tasks, highlighting the need for productivity. In this section, we will assess productivity by evaluating the validity of LLM-generated examples based on given example pairs from ARC tasks.
While productivity ideally involves testing for infinite generative capacity, practical limitations necessitate alternative approaches. The challenge lies in demonstrating that a system can produce an unlimited number of novel, meaningful outputs from a finite set of inputs and rules. Previous studies have addressed this challenge by examining whether valid outputs can be produced under added constraints (van der Velde et al., 2004; Lake and Baroni, 2018; Hupkes et al., 2020). These constraints serve to create a more manageable testing environment while still allowing for the assessment of generative capabilities. Following this methodology, our study investigates how effectively LLMs can generate valid outputs when presented with an ARC task and its underlying conceptual rule. This approach allows us to evaluate productivity within a controlled framework while still capturing the essence of generative capacity.
To understand how well LLMs can generate new expressions based on inherent logical concepts, we conduct experiments using ARC tasks. Productivity in this context involves two main steps: 1) inferring specific rules for image generation from example images and natural language expressions, and 2) applying these rules to generate new, unseen images. However, the standard approach to solving ARC tasks, as explored in previous sections, is insufficient to confirm these two processes. Therefore, we propose a novel experiment: Given an ARC task and a basic rule shared with similar ARC tasks, can LLMs generate valid examples of the given task? If LLMs can understand the relationship between the given ARC task and the abstract rule, they should be able to infer specific rules for the task and generate new valid examples. Through this, we aim to determine whether LLMs can imitate the productivity of human thinking in generating novel solutions.
3.3.2. Validity of Augmentation
To precisely evaluate whether LLMs can infer their own generation rules given ARC examples and create new tasks by appropriately applying these rules, we rigorously controlled the prompts. LLMs receive two types of prompts: example pairs included in the ARC task and descriptions of abstract rules applicable to similar ARC tasks. However, in this case, one example pair was used as the basis for generation, and the remaining examples were used for inferring specific rules for the task. Based on the category of ConceptARC (Moskvichev et al., 2023), which organizes a subset of ARC tasks into 16 distinct categories according to human classification criteria, we developed abstract rules. For each category within ConceptARC, we crafted a corresponding abstract rule, ensuring that tasks within the same category adhere to the identical abstract rule. An example of this abstract rule is shown at the top of the Inverse Transformation Prompt panel in Fig. 13.
We proposed the ITP, a prompting technique for this experiment. ITP instructs LLMs to generate multiple valid examples using the ARC task and its abstract rules. Fig. 13 demonstrates the process by which LLMs generate new examples given the ARC task and the corresponding ITP. LLMs generate multiple inputs that can form pairs with the output from one example of the task. This example, as mentioned earlier, is selected to serve as the basis for generation and is not included in ITP. If LLMs understood the specific rules corresponding to the ARC task given through ITP, the new example pairs generated by LLMs would be suitable as examples of the task.
ITP is based on many-to-one corresponding to elicit two advantages. First, the method of generating only input is more data-efficient than the method of generating both input and output because the output of the existing ARC task can be used as is. Since all tasks in ARC have example pairs, reusing these examples can be said to make full use of the given data. ITP allows for the reuse of a single ARC task multiple times, making it data-efficient. In particular, using ITP can further increase data efficiency by allowing one ARC task to be reused multiple times by changing the order of examples. Secondly, ITP increases the likelihood of generating valid responses. Through simple simulations, we have seen that inferring inputs from output tends to be more likely to generate valid results than inferring outputs from input. Because generating input from output is subject to relatively less stringent constraints, there is often a wide range of acceptable outcomes.
In the process of creating ITP, we encounter two challenges. First, according to the ConceptARC category, there could be multiple solutions within one category. Fig. 14(a) illustrates that there are various types of tasks with the same category. Abstract rules given in the same sentences for each category may not be sufficient to cover various types of tasks. Second, there were ARC tasks that made not possible to infer multiple inputs from a single output (Fig. 14(b)). In such cases, there was only one valid input. Although we tried to take these cases into account while writing the ITP, these challenges nevertheless harmed the experimental results.
Before analyzing the experimental results, it was necessary to redefine the evaluation metric because the focus shifted from solving ARC tasks to generating valid examples. As previously explained, for a given example of a particular ARC task, we generated valid inputs that could be paired with the corresponding output. To successfully generate these inputs, the LLM must understand the specific rules of the given ARC task through its ITP and apply those understood rules to the output to create valid inputs. In this experiment, we evaluated whether all generated inputs were valid for each ARC task. This evaluation metric assesses both the LLM’s understanding of the correct rules and its ability to generate valid examples based on those rules. Consequently, this experiment systematically evaluates the LLMs’ capability to generate logical and valid demo pairs, enhancing our understanding of their ability to create new representations.
| Category | Generated Examples | Valid Examples | Validity |
| Above Below | 158 | 34 | 21.52% |
| Center | 236 | 35 | 14.83% |
| Clean Up | 183 | 83 | 45.36% |
| Complete Shape | 147 | 37 | 25.17% |
| Copy | 153 | 4 | 2.61% |
| Count | 202 | 29 | 14.36% |
| Extend To Boundary | 167 | 8 | 4.79% |
| Extract Objects | 176 | 21 | 11.93% |
| Filled Not Filled | 203 | 29 | 14.29% |
| Horizontal Vertical | 114 | 7 | 6.14% |
| Inside Outside | 191 | 24 | 12.57% |
| Move To Boundary | 165 | 12 | 7.27% |
| Order | 162 | 26 | 16.05% |
| Same Different | 246 | 76 | 30.89% |
| Top Bottom 2D | 255 | 59 | 23.14% |
| Top Bottom 3D | 215 | 25 | 11.63% |
| Total | 2,913 | 509 | 17.12% |
Based on 160 ARC tasks classified by ConceptARC, we evaluated the validity of a total of 2,913 generated examples. The average valid generation ratio was approximately 17.1%, with the remaining examples deemed invalid. As previously mentioned, the validity of these results was determined by human judgment, assessing whether the generated tasks adhered to the given rules. The results in Table LABEL:tbl:productivity/generation_validity show that LLMs exhibit some capability in generating examples that align with the specified rules. However, due to weak criteria for determining validity, there is a limitation: even if an infinite number of results can be generated, they cannot be reliably used without post-processing the data.
3.3.3. Case Study: Invalid Production
An analysis of the generated inputs was conducted to investigate the reasons behind LLMs’ inability to produce valid inputs for ARC tasks. Two major limitations were observed when LLMs generated new ARC tasks. First, LLMs tended to simply copy inputs rather than infer meaningful rules from given example pairs. As shown in Fig. 15, this occurred repeatedly despite attempts to prevent it through prompts. Second, LLMs failed to properly consider the steps needed to generate inputs from given outputs. This frequently resulted in the creation of examples that could not be solved by the specific rules of the task. For instance, in cases where all vertices of a square were erased in the input, it became impossible to determine the color of the vertices, making it infeasible to infer the given output. These limitations suggest that LLMs lack an understanding of the semantics applicable in ARC tasks and the ability to compose these semantics according to constraints.
3.3.4. Conclusion
In Section 3.3, we conducted experiments to confirm the productivity of LLMs by assessing whether they can understand given ARC tasks in abstracted representations and generate valid new examples based on abstracted rules. Although it is known that LLMs have great strengths in creating creative works, our experimental results reveal that LLMs are weak in understanding rules and producing creations that adhere to those rules. Moreover, the observed limitations highlight a critical gap in LLMs’ ability to engage in higher-level reasoning and abstraction, which are essential for successfully solving ARC tasks that require an understanding of underlying principles rather than surface patterns. These results suggest that when LLMs generate outputs, they tend to mimic human-created results rather than truly understanding and applying rules. This makes it difficult for LLMs to reach the level of generation that humans can achieve.
Similarly, previous studies have shown similar results in measuring the productivity of AI models. Researchers tested how well pre-LLM models generalize to novel command combinations (Lake and Baroni, 2018; van der Velde et al., 2004). Their findings revealed strong performance on trained data but weaknesses in generating responses to unseen commands. Some researchers argued that LLMs struggle with generation under complex constraints and proposed improved models to address this issue (Koh et al., 2024; Liu et al., 2024). They propose novel frameworks to enhance LLMs for generating desired outputs when complex constraints are introduced, rather than relying solely on the base models. These researches share similarities with our study, which encountered difficulties in augmenting valid ARC tasks based on complex rules.
4. Discussion
Through the three experiments in Section 3, we have observed that LLMs demonstrate strengths in understanding and manipulating both image and text inputs. However, they still exhibit weaknesses in logical inference, sequential planning based on understanding, and generating unseen images according to predefined rules. We will conclude by introducing the current research directions aimed at further enhancing LLMs’ ability and outlining the goals after solving ARC.
4.1. What Should LLMs Possess to Solve ARC?
Based on the experimental results of Section 3, it is evident that LLMs still cannot solve ARC effectively. This is attributed to the deficiencies in logical coherence, compositionality, and productivity. How can we improve the inference capabilities of LLMs? In this section, we explore directions to enhance LLMs from the perspectives of abstraction knowledge and reasoning.
4.1.1. Abstract Knowledge
To solve ARC, the first challenge is the ability to extract the implicit information contained within ARC. Xu et al. argued that object-based representation is crucial for solving ARC and proposed ARGA (Xu et al., 2023a), which converts given example grids into graphs. Their follow-up study (Xu et al., 2023b) involved LLM solving ARC tasks using information obtained from ARGA and showed notable performance for object-based ARC tasks. However, these studies have a fundamental weakness in that they cannot be applied to ARC tasks without objects. Since only about 40% of ARC tasks contain object concepts (Xu et al., 2023a), this method cannot be applied to more than half of the tasks. Wang et al. on the other hand, improved the abstraction ability of LLMs to some extent with a graph-form dataset consisting of 221K textual descriptions, called AbsPyramid (Wang et al., 2023b), and also proposed a framework called AbsInstruct (Wang et al., 2024a) utilizing this dataset. Attempting to structure sentences can be an effective abstraction method for natural language, but its effectiveness cannot be seen in tasks that do not contain sentences.
4.1.2. Reasoning
Another challenge for LLMs in the context of ARC is the vast search space. One method gaining attention to address this is to enable LLMs to generate DSLs themselves. Rajani et al. introduced CAGE (Rajani et al., 2019), which prompts LLMs to generate explanations before generating answers. Subsequently, Wang et al. (Wang et al., 2024b) reported improved results by having LLMs generate DSLs based on hypotheses they set themselves. Additionally, active research is underway on prompting techniques applying algorithmic approaches. Zhou et al. (Zhou et al., 2022) demonstrated enhanced inference performance in LLMs by applying in-context learning. Follow-up research is actively being conducted following CoT and ToT. For example, CoT-SC (Wang et al., 2023c) is a study that selects results through voting from multiple instances of CoT, GoT (Besta et al., 2023) secures flexibility by enabling the generation of graph-like thought nodes, and XoT (Ding et al., 2023) uses the thought tree while Monte Carlo tree search and refines the tree with reinforcement learning. However, these attempts are closer to additional learning for LLMs, and more researches are needed to ascertain whether fundamental improvements in LLMs’ reasoning abilities are achievable.
4.2. Limitation of ARC
Does solving ARC signify the completion of human-like AI? To answer this question, two doubts need to be appropriately addressed: 1) Will the ARC solver possess human-level problem-solving abilities? and 2) Will that solver think like humans to solve ARC? It’s not easy to imagine how the ARC solver operates without human-level reasoning. At this point, what we can assume is that the model will have the three properties of LoTH, and the model could be capable of several types of reasoning included in ARC. With this hypothesis, we attempt to address the following questions.
4.2.1. Will the Model Possess Human-Level Problem-Solving Abilities?
Being capable of reasoning does not necessarily equate to having human-level problem-solving abilities. In other words, even if a model can reason to a level that can solve ARC, it may not have human-level problem-solving capabilities. Various tasks that humans face are generally more complex than ARC and involve various other cognitive factors besides reasoning. Therefore, even models that can solve ARC may have the following limitations compared to human-level problem-solving abilities.
First, with the current ARC criteria, it’s still unknown whether the model that solved it can solve more complex types of tasks. This is because ARC tasks focus on just reasoning and are therefore presented in a relatively simple environment. Whether the reasoning ability learned through ARC would also work in more complex environments has not been revealed. Second, solving ARC does not imply the presence of other components of intelligence beyond reasoning. While reasoning is undoubtedly a core aspect of cognitive processes, it is not the entirety of intelligence. There is research shows that solving human-level complex tasks requires various cognitive abilities (Gardner, 2011).
4.2.2. Will the Model Think Like Humans?
Even if we assume that the ARC solver can reason in terms of LoTH, we cannot guarantee whether this solver’s process is human-like for the following two reasons. Firstly, the current ARC provides a performance measure that rewards only for solving a task. It’s important to recognize that such a measure might instigate a wrong purpose, leading to what is known as the King Midas problem (Russell and Norvig, 1995). This problem emphasizes the risk of AI achieving its given objective too literally, leading to unintended negative consequences, underscoring the importance of aligning AI’s goals with human values and the broader context. The policy of rewarding only the results, excluding the solution process, makes it difficult to evaluate whether the solution process is similar to human reasoning. Therefore, models trained on current ARC likely differ in how they solve tasks compared to humans. The second reason is that directly comparing the reasoning processes of humans and language models is challenging. The process by which humans solve ARC tasks has not been investigated, making it unclear how the solving process differs between humans and artificial intelligence. Furthermore, there is a lack of metrics for comparing the solving processes, making direct comparisons difficult.
4.3. Future Direction After Solving ARC
To summarize, solving ARC tasks does not directly imply achieving human-level artificial intelligence. Moreover, there is a challenge in comparing task-solving approaches with those of humans. Thus, we suggest three alternatives to more accurately measure human-level inference abilities.
4.3.1. Using Different Benchmarks
One limitation of ARC is its simple environment. SQA3D (Ma et al., 2023), for instance, addresses inference tasks in a 3D domain by extending them into question-answering tasks using simulators like ScanNet (Dai et al., 2017). Additionally, benchmarks such as TGIF-QA (Jang et al., 2017), MovieQA (Tapaswi et al., 2016), TVQA (Lei et al., 2018), and STAR (Wu et al., 2021), which append question-answering to videos, have been proposed. Such benchmarks mimicking real-world inference scenarios could serve as supplements to measure complex abstractions not covered by ARC.
4.3.2. Quantification of ARC Task-Solving Processes
Chollet, the creator of ARC, argued that ARC maximizes generality while minimizing prior and experience (Chollet, 2019), but these components have not been quantitatively evaluated. As a result, the quantitative assessment of factors such as the generality achieved by models solving ARC, the level of prior knowledge, and the components of prior knowledge remains elusive. One possible way to quantitatively evaluate the process of solving ARC tasks is to quantify the model’s achievement of prior, experience, and generality.
4.3.3. Adding Evaluation Methods to Compare Task-Solving Processes with Human Approaches
Recent ARC research has focused on finding ways for AI to solve tasks. However, there are doubts about how similar these solutions are to those of humans. The initial paper by Johnson et al. (Johnson et al., 2021) analyzed human ARC solutions. Subsequently, LARC (Acquaviva et al., 2022) was proposed to analyze how tasks are solved through the language-based explanation of human solutions. Tools for facilitating the collection of human data are also continuously being developed. Kim et al. (Kim et al., 2022), for instance, have analyzed how tasks are solved through O2ARC. It is suggested to not only calculate simple correctness for each ARC task but also to add similarity with human data to the evaluation.
4.4. Recent Research Trends on the Reasoning Abilities of LLMs
In this paper, we utilized the ARC to evaluate and enhance the reasoning capabilities of LLMs. ARC serves as a crucial benchmark for testing AI models’ ability to perform human-like reasoning, and it plays a central role in our research. Beyond ARC, various other datasets can be leveraged to strengthen LLMs’ reasoning abilities further. For instance, the DROP (Discrete Reasoning Over Paragraphs) dataset (Dua et al., 2019) evaluates a model’s capability to perform logical reasoning over complex text and synthesize information from multiple sources to generate accurate answers. Similarly, CommonsenseQA (Talmor et al., 2019) supports training models to perform commonsense reasoning, enabling them to make logical decisions in everyday scenarios. The BoolQ dataset (Clark et al., 2019) requires models to use logical reasoning to answer true/false questions, while the GSM8K dataset (Cobbe et al., 2021) focuses on enhancing step-by-step logical and mathematical reasoning skills through high-quality problem-solving tasks. Together, these datasets, alongside ARC, provide invaluable resources for systematically enhancing the diverse reasoning capabilities of LLMs.
However, recent studies indicate that despite their proficiency in language-based tasks, LLMs still exhibit significant limitations in their reasoning abilities. Carvalho et al. (de Carvalho et al., 2024) explored the reasoning capabilities of models like GPT-3.5 and GPT-4 in non-linguistic tasks requiring strategic thinking and spatial reasoning tasks. They found that while LLMs perform well on basic language tasks, they struggle with reasoning and decision-making in tasks beyond their training data, demonstrating limited cognitive flexibility and generalization. Similarly, Gendron et al. (Gendron et al., 2024) critically examined the abstract reasoning skills of LLMs, revealing poor performance on tasks requiring the identification and application of general patterns from only a small number of examples, specifically designed to test broad generalization abilities. These studies collectively highlight that current LLMs, though advanced in linguistic tasks, are still far from achieving robust reasoning abilities across diverse domains.
To build on this foundation, several advanced approaches have been developed to further enhance LLMs’ reasoning capabilities. Reinforcement learning with human feedback (Christiano et al., 2017) allows models to learn and refine their reasoning processes through feedback during training. The Chain of Thought (CoT) prompting technique (Wei et al., 2022) facilitates multi-step reasoning by breaking down complex tasks into smaller, manageable steps. Reasoning-centric fine-tuning (Lewkowycz et al., 2022) focuses on improving performance in specific reasoning tasks by leveraging specialized datasets. Additionally, incorporating knowledge graphs during pre-training (Liu et al., 2020) enhances logical inference by integrating structured world knowledge. Finally, explainable AI techniques, such as self-reflective learning, empower LLMs to identify and correct their reasoning errors, thereby contributing to greater transparency and trustworthiness in AI (Biran and Cotton, 2017). Collectively, these approaches play a crucial role in advancing LLMs’ reasoning capabilities across various domains.
Moreover, recent research has introduced innovative approaches to further augment the reasoning capabilities of LLMs. For example, multimodal learning techniques allow LLMs to simultaneously process text and image data, thereby enhancing their ability to tackle complex reasoning tasks (Singh et al., 2022). Adaptive learning strategies that incorporate human feedback have also been shown to improve the models’ adaptability across a wide range of tasks (Ouyang et al., 2022). Additionally, integrating programming languages with LLMs has been proposed as a means to enhance logical problem-solving abilities (Gao et al., 2023). These cutting-edge studies significantly contribute to systematically strengthening the multidimensional reasoning capabilities of LLMs.
5. Conclusions
In this study, we addressed the limitation of existing research, which predominantly analyzed LLM’s inference ability from a deterministic perspective, by introducing the LoTH perspective to ensure a fair evaluation of the process as well. By comparing and analyzing arguments across various domains of inference, we confirmed that logical coherence, compositionality, and productivity are quantifiable components to evaluate inference ability. Next, we proposed three experiments using the ARC to quantitatively evaluate these three components from the LoTH perspective. The results showed that although current LLMs exhibit outstanding performance, they lack logical coherence, compositionality, and productivity in their processes, suggesting that they are closer to probabilistic mimicry rather than possessing autonomous reasoning abilities. Finally, we explored meaningful research directions for LLM to acquire reasoning abilities from the LoTH perspective, as well as alternative approaches beyond ARC. This attempt to quantitatively evaluate the inference process from the LoTH perspective through experimental methods represents a differentiated approach not present in previous research. It contributes by providing a new perspective on how to handle reasoning abilities in the field of computer science, going beyond LLMs.
Acknowledgements.
This work was supported by the IITP (RS-2023-00216011, RS-2024-00445087, No. 2019-0-01842) and the NRF (RS-2024-00451162) grants funded by the Ministry of Science and ICT, Korea. Experiments are supported by the Accelerate Foundation Models Research Initiative, Microsoft.References
- (1)
- Acquaviva et al. (2022) Samuel Acquaviva, Yewen Pu, Marta Kryven, Theodoros Sechopoulos, Catherine Wong, Gabrielle Ecanow, Maxwell Nye, Michael Tessler, and Joshua B. Tenenbaum. 2022. Communicating Natural Programs to Humans and Machines. In NeurIPS.
- Allen-Zhu and Li (2023) Zeyuan Allen-Zhu and Yuanzhi Li. 2023. Physics of Language Models: Part 3.2, Knowledge Manipulation. arXiv:2309:14402 (2023).
- Barrett et al. (2018) David Barrett, Felix Hill, Adam Santoro, Ari Morcos, and Timothy Lillicrap. 2018. Measuring Abstract Reasoning in Neural Networks. In ICML.
- Besta et al. (2023) Maciej Besta, Nils Blach, Ales Kubicek, Robert Gerstenberger, Lukas Gianinazzi, Joanna Gajda, Tomasz Lehmann, Michal Podstawski, Hubert Niewiadomski, Piotr Nyczyk, and Torsten Hoefler. 2023. Graph of Thoughts: Solving Elaborate Problems with Large Language Models. arXiv:2308.09687 (2023).
- Biran and Cotton (2017) Or Biran and Courtenay Cotton. 2017. Explanation and Justification in Machine Learning: A Survey. In IJCAI Workshop.
- Borsky (2021) Alexey Borsky. 2021. ARC-Game. https://github.com/volotat/ARC-Game
- Bubeck et al. (2023) Sébastien Bubeck, Varun Chandrasekaran, Ronen Eldan, Johannes Gehrke, Eric Horvitz, Ece Kamar, Peter Lee, Yin Tat Lee, Yuanzhi Li, Scott Lundberg, Harsha Nori, Hamid Palangi, Marco Tulio Ribeiro, and Yi Zhang. 2023. Sparks of Artificial General Intelligence: Early Experiments with GPT-4. arXiv:2303.12712 (2023).
- Chang et al. (2024) Yupeng Chang, Xu Wang, Jindong Wang, Yuan Wu, Linyi Yang, Kaijie Zhu, Hao Chen, Xiaoyuan Yi, Cunxiang Wang, Yidong Wang, Wei Ye, Yue Zhang, Yi Chang, Philip S. Yu, Qiang Yang, and Xing Xie. 2024. A Survey on Evaluation of Large Language Models. ACM Transactions on Intelligent Systems and Technology (2024).
- Chollet (2019) François Chollet. 2019. On the Measure of Intelligence. arXiv:1911.01547 (2019).
- Christiano et al. (2017) Paul F Christiano, Jan Leike, Tom Brown, Miljan Martic, Shane Legg, and Dario Amodei. 2017. Deep Reinforcement Learning from Human Preferences. In NeurIPS.
- Clark et al. (2019) Christopher Clark, Kenton Lee, Ming-Wei Chang, Tom Kwiatkowski, Michael Collins, and Kristina Toutanova. 2019. BoolQ: Exploring the Surprising Difficulty of Natural Yes/No Questions. In NAACL.
- Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. 2021. Training Verifiers to Solve Math Word Problems. arXiv:2110.14168 (2021).
- Dai et al. (2017) Angela Dai, Angel X Chang, Manolis Savva, Maciej Halber, Thomas Funkhouser, and Matthias Nießner. 2017. ScanNet: Richly-Annotated 3D Reconstructions of Indoor Scenes. In CVPR. 5828–5839.
- de Carvalho et al. (2024) Gonçalo Hora de Carvalho, Robert Pollice, and Oscar Knap. 2024. Show, Don’t Tell: Evaluating Large Language Models Beyond Textual Understanding with ChildPlay. arXiv:2407.11068 (2024).
- Ding et al. (2023) Ruomeng Ding, Chaoyun Zhang, Lu Wang, Yong Xu, Minghua Ma, Wei Zhang, Si Qin, Saravan Rajmohan, Qingwei Lin, and Dongmei Zhang. 2023. Everything of Thoughts: Defying the Law of Penrose Triangle for Thought Generation. arXiv:2311.04254 (2023).
- Dua et al. (2019) Dheeru Dua, Yizhong Wang, Pradeep Dasigi, Gabriel Stanovsky, Sameer Singh, and Matt Gardner. 2019. DROP: A Reading Comprehension Benchmark Requiring Discrete Reasoning Over Paragraphs. In NAACL.
- Dziri et al. (2023) Nouha Dziri, Ximing Lu, Melanie Sclar, Xiang Lorraine Li, Liwei Jiang, Bill Yuchen Lin, Peter West, Chandra Bhagavatula, Ronan Le Bras, Jena D. Hwang, Soumya Sanyal, Sean Welleck, Xiang Ren, Allyson Ettinger, Zaid Harchaoui, and Yejin Choi. 2023. Faith and Fate: Limits of Transformers on Compositionality. In NeurIPS.
- Fodor (1975) Jerry A Fodor. 1975. The Language of Thought. Vol. 5. Harvard University Press.
- Fodor and Pylyshyn (1988) Jerry A Fodor and Zenon W Pylyshyn. 1988. Connectionism and Cognitive Architecture: A Critical Analysis. Cognition 28, 1-2 (1988), 3–71.
- Gao et al. (2023) Luyu Gao, Aman Madaan, Shuyan Zhou, Uri Alon, Pengfei Liu, Yiming Yang, Jamie Callan, and Graham Neubig. 2023. PAL: Program-Aided Language Models. In ICML. 10764–10799.
- Gardner (2011) Howard E Gardner. 2011. Frames of Mind: The Theory of Multiple Intelligences. Basic Books.
- Gendron et al. (2024) Gaël Gendron, Qiming Bao, Michael Witbrock, and Gillian Dobbie. 2024. Large Language Models are Not Strong Abstract Reasoners. IJCAI (2024).
- Gurnee and Tegmark (2023) Wes Gurnee and Max Tegmark. 2023. Language Models Represent Space and Time. arXiv:2310.02207 (2023).
- Hill et al. (2019) Felix Hill, Adam Santoro, David GT Barrett, Ari S Morcos, and Timothy Lillicrap. 2019. Learning to Make Analogies by Contrasting Abstract Relational Structure. In ICLR.
- Hodel (2024) Michael Hodel. 2024. Addressing the Abstraction and Reasoning Corpus via Procedural Example Generation. arXiv:2404.07353 (2024). https://github.com/michaelhodel/re-arc?tab=readme-ov-file
- Huang and Chang (2023) Jie Huang and Kevin Chen-Chuan Chang. 2023. Towards Reasoning in Large Language Models: A Survey. In ACL Findings.
- Hupkes et al. (2020) Dieuwke Hupkes, Verna Dankers, Mathijs Mul, and Elia Bruni. 2020. Compositionality Decomposed: How Do Neural Networks Generalise? Journal of Artificial Intelligence Research 67 (2020), 757–795.
- Jang et al. (2017) Yunseok Jang, Yale Song, Youngjae Yu, Youngjin Kim, and Gunhee Kim. 2017. TGIF-QA: Toward Spatio-Temporal Reasoning in Visual Question Answering. In CVPR. 2758–2766.
- Johnson et al. (2021) Aysja Johnson, Wai Keen Vong, Brenden M. Lake, and Todd M. Gureckis. 2021. Fast and Flexible: Human Program Induction in Abstract Reasoning Tasks. In CogSci.
- Kim et al. (2022) Subin Kim, Prin Phunyaphibarn, Donghyun Ahn, and Sundong Kim. 2022. Playgrounds for Abstraction and Reasoning. In NeurIPS Workshop on nCSI.
- Koh et al. (2024) Jing Yu Koh, Daniel Fried, and Russ R Salakhutdinov. 2024. Generating Images with Multimodal Language Models. In NeurIPS, Vol. 36.
- Lab42 (2024) Lab42. 2024. ARC-PRIZE competition. https://arcprize.org/
- Lake and Baroni (2018) Brenden Lake and Marco Baroni. 2018. Generalization without Systematicity: On the Compositional Skills of Sequence-to-Sequence Recurrent Networks. In ICML. PMLR, 2873–2882.
- Lei et al. (2018) Jie Lei, Licheng Yu, Mohit Bansal, and Tamara L Berg. 2018. TVQA: Localized, Compositional Video Question Answering. In EMNLP.
- Lewkowycz et al. (2022) Aitor Lewkowycz, Anders Andreassen, David Dohan, Ethan Dyer, Henryk Michalewski, Vinay Ramasesh, Ambrose Slone, Cem Anil, Imanol Schlag, Theo Gutman-Solo, Yuhuai Wu, Behnam Neyshabur, Guy Gur-Ari, and Vedant Misra. 2022. Solving Quantitative Reasoning Problems with Language Models. In NeurIPS.
- Li et al. (2024) Zhaoyi Li, Gangwei Jiang, Hong Xie, Linqi Song, Defu Lian, and Ying Wei. 2024. Understanding and Patching Compositional Reasoning in LLMs. arXiv:2402.14328 (2024).
- Liu et al. (2024) Mushui Liu, Yuhang Ma, Xinfeng Zhang, Yang Zhen, Zeng Zhao, Zhipeng Hu, Bai Liu, and Changjie Fan. 2024. LLM4GEN: Leveraging Semantic Representation of LLMs for Text-to-Image Generation. arXiv:2407.00737 (2024).
- Liu et al. (2020) Weijie Liu, Peng Zhou, Zhe Zhao, Zhiruo Wang, Qi Ju, Haotang Deng, and Ping Wang. 2020. K-BERT: Enabling Language Representation with Knowledge Graphs. In AAAI.
- Lu et al. (2024) Pan Lu, Hritik Bansal, Tony Xia, Jiacheng Liu, Chunyuan Li, Hannaneh Hajishirzi, Hao Cheng, Kai-Wei Chang, Michel Galley, and Jianfeng Gao. 2024. MathVista: Evaluating Math Reasoning in Visual Contexts with GPT-4V, Bard, and Other Large Multimodal Models. In ICLR.
- Ma et al. (2024) Teli Ma, Rong Li, and Junwei Liang. 2024. An Examination of the Compositionality of Large Generative Vision-Language Models. In NAACL.
- Ma et al. (2023) Xiaojian Ma, Silong Yong, Zilong Zheng, Qing Li, Yitao Liang, Song-Chun Zhu, and Siyuan Huang. 2023. SQA3D: Situated Question Answering in 3D Scenes. In ICLR.
- Mańdziuk and Żychowski (2019) Jacek Mańdziuk and Adam Żychowski. 2019. DeepIQ: A Human-Inspired AI System for Solving IQ Test Problems. In IJCNN.
- Matzen et al. (2010) Laura E Matzen, Zachary O Benz, Kevin R Dixon, Jamie Posey, James K Kroger, and Ann E Speed. 2010. Recreating Raven’s: Software for Systematically Generating Large Numbers of Raven-Like Matrix Problems With Normed Properties. Behavior Research Methods 42, 2 (2010), 525–541.
- McCulloch and Pitts (1943) Warren S McCulloch and Walter Pitts. 1943. A Logical Calculus of the Ideas Immanent in Nervous Activity. The Bulletin of Mathematical Biophysics 5 (1943), 115–133.
- Mikołaj and Jacek (2023) Małkiński Mikołaj and Mańd́ziuk Jacek. 2023. A Review of Emerging Research Directions in Abstract Visual Reasoning. Information Fusion 91 (2023), 713–736.
- Mirchandani et al. (2023) Suvir Mirchandani, Fei Xia, Pete Florence, Brian Ichter, Danny Driess, Montserrat Gonzalez Arenas, Kanishka Rao, Dorsa Sadigh, and Andy Zeng. 2023. Large Language Models as General Pattern Machines. arXiv:2307.04721 (2023).
- Moskvichev et al. (2023) Arseny Moskvichev, Victor Vikram Odouard, and Melanie Mitchell. 2023. The ConceptARC Benchmark: Evaluating Understanding and Generalization in the ARC Domain. Transactions on Machine Learning Research (2023).
- Nie et al. (2020) Weili Nie, Zhiding Yu, Lei Mao, Ankit B Patel, Yuke Zhu, and Anima Anandkumar. 2020. Bongard-Logo: A New Benchmark for Human-Level Concept Learning and Reasoning. In NeurIPS.
- Ouyang et al. (2022) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simense, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022. Training Language Models to Follow Instructions with Human Feedback. In NeurIPS.
- Park et al. (2023) Jaehyun Park, Jaegyun Im, Sanha Hwang, Mintaek Lim, Sabina Ualibekova, Sejin Kim, and Sundong Kim. 2023. Unraveling the ARC Puzzle: Mimicking Human Solutions with Object-Centric Decision Transformer. ICML Workshop on ILHF (2023).
- Qiu et al. (2024) Linlu Qiu, Liwei Jiang, Ximing Lu, Melanie Sclar, Valentina Pyatkin, Chandra Bhagavatula, Bailin Wang, Yoon Kim, Yejin Choi, Nouha Dziri, and Xiang Ren. 2024. Phenomenal Yet Puzzling: Testing Inductive Reasoning Capabilities of Language Models with Hypothesis Refinement. In ICLR.
- Rajani et al. (2019) Nazneen Fatema Rajani, Bryan McCann, Caiming Xiong, and Richard Socher. 2019. Explain Yourself! Leveraging Language Models for Commonsense Reasoning. In ACL.
- Rosenblatt (1958) Frank Rosenblatt. 1958. The Perceptron: A Probabilistic Model for Information Storage and Organization in the Brain. Psychological Review 65, 6 (1958), 386–408.
- Russell and Norvig (1995) Stuart J Russell and Peter Norvig. 1995. Artificial Intelligence: A Modern Approach. Pearson.
- Shin et al. (2023) Donghyeon Shin, Sanha Hwang, Seokki Lee, Yunho Kim, and Sundong Kim. 2023. MC-LARC Benchmark to Measure LLM Reasoning Capability. In Korea Software Congress.
- Singh et al. (2022) Amanpreet Singh, Ronghang Hu, Vedanuj Goswami, Guillaume Couairon, Wojciech Galuba, Marcus Rohrbach, and Douwe Kiela. 2022. FLAVA: A Foundational Language and Vision Alignment Model. In CVPR.
- Sinha et al. (2024) Sania Sinha, Tanawan Premsri, and Parisa Kordjamshidi. 2024. A Survey on Compositional Learning of AI Models: Theoretical and Experimental Practices. arXiv:2406.08787 (2024).
- Talmor et al. (2019) Alon Talmor, Jonathan Herzig, Nicholas Lourie, and Jonathan Berant. 2019. CommonsenseQA: A Question Answering Challenge Targeting Commonsense Knowledge. In NAACL.
- Tapaswi et al. (2016) Makarand Tapaswi, Yukun Zhu, Rainer Stiefelhagen, Antonio Torralba, Raquel Urtasun, and Sanja Fidler. 2016. MovieQA: Understanding Stories in Movies through Question-Answering. In CVPR. 4631–4640.
- Turing (1950) Alan Turing. 1950. Computing Machinery and Intelligence. Mind 59, 236 (1950), 433–460.
- Tyen et al. (2024) Gladys Tyen, Hassan Mansoor, Victor Cărbune, Yuanzhu Peter Chen, and Tony Mak. 2024. LLMs cannot find reasoning errors, but can correct them given the error location. In ACL Findings.
- Valmeekam et al. (2024) Karthik Valmeekam, Alberto Olmo, Sarath Sreedharan, and Subbarao Kambhampati. 2024. PlanBench: An Extensible Benchmark for Evaluating Large Language Models on Planning and Reasoning about Change. In NeurIPS.
- van der Velde et al. (2004) Frank van der Velde, Gwendid T van der Voort van der Kleij, and Marc de Kamps. 2004. Lack of Combinatorial Productivity in Language Processing with Simple Recurrent Networks. Connection Science 16, 1 (2004), 21–46.
- Wang et al. (2023a) Boshi Wang, Sewon Min, Xiang Deng, Jiaming Shen, You Wu, Luke Zettlemoyer, and Huan Sun. 2023a. Towards Understanding Chain-of-Thought Prompting: An Empirical Study of What Matters. In ACL.
- Wang et al. (2019) Haoyu Wang, Mo Yu, Xiaoxiao Guo, Rajarshi Das, Wenhan Xiong, and Tian Gao. 2019. Do Multi-Hop Readers Dream of Reasoning Chains?. In EMNLP Workshop on MRQA.
- Wang and Su (2015) Ke Wang and Zhendong Su. 2015. Automatic Generation of Raven’s Progressive Matrices. In IJCAI.
- Wang et al. (2024b) Ruocheng Wang, Eric Zelikman, Gabriel Poesia, Yewen Pu, Nick Haber, and Noah D Goodman. 2024b. Hypothesis Search: Inductive Reasoning with Language Models. In ICLR.
- Wang et al. (2023c) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc Le, Ed Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou. 2023c. Self-Consistency Improves Chain of Thought Reasoning in Language Models. In ICLR.
- Wang et al. (2024a) Zhaowei Wang, Wei Fan, Qing Zong, Hongming Zhang, Sehyun Choi, Tianqing Fang, Xin Liu, Yangqiu Song, Ginny Y Wong, and Simon See. 2024a. AbsInstruct: Eliciting Abstraction Ability from LLMs through Explanation Tuning with Plausibility Estimation. arXiv:2402.10646 (2024).
- Wang et al. (2023b) Zhaowei Wang, Haochen Shi, Weiqi Wang, Tianqing Fang, Hongming Zhang, Sehyun Choi, Xin Liu, and Yangqiu Song. 2023b. AbsPyramid: Benchmarking the Abstraction Ability of Language Models with a Unified Entailment Graph. arXiv:2311.09174 (2023).
- Webb et al. (2020) Taylor Webb, Zachary Dulberg, Steven Frankland, Alexander Petrov, Randall O’Reilly, and Jonathan Cohen. 2020. Learning Representations That Support Extrapolation. In ICML.
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, and Denny Zhou. 2022. Chain-of-Thought Prompting Elicits Reasoning in Large Language Models. In NeurIPS.
- West et al. (2024) Peter West, Ximing Lu, Nouha Dziri, Faeze Brahman, Linjie Li, Jena D Hwang, Liwei Jiang, Jillian Fisher, Abhilasha Ravichander, Khyathi Chandu, Benjamin Newman, Pang Wei Koh, Allyson Ettinger, and Yejin Choi. 2024. The Generative AI Paradox: “What It Can Create, It May Not Understand”. In ICLR.
- Wiener (1950) Norbert Wiener. 1950. Cybernetics. Bulletin of the American Academy of Arts and Sciences 3, 7 (1950), 2–4.
- Wind (2020) Johan Sokrates Wind. 2020. ARC-Solution. https://github.com/top-quarks/ARC-solution
- Wu et al. (2021) Bo Wu, Shoubin Yu, Zhenfang Chen, Joshua B Tenenbaum, and Chuang Gan. 2021. STAR: A Benchmark for Situated Reasoning in Real-World Videos. In NeurIPS.
- Xu et al. (2023a) Yudong Xu, Elias B. Khalil, and Scott Sanner. 2023a. Graphs, Constraints, and Search for the Abstraction and Reasoning Corpus. In AAAI.
- Xu et al. (2023b) Yudong Xu, Wenhao Li, Pashootan Vaezipoor, Scott Sanner, and Elias B Khalil. 2023b. LLMs and the Abstraction and Reasoning Corpus: Successes, Failures, and the Importance of Object-based Representations. Transactions on Machine Learning Research (2023).
- Xue et al. (2023) Cheng Xue, Vimukthini Pinto, Chathura Gamage, Ekaterina Nikonova, Peng Zhang, and Jochen Renz. 2023. Phy-Q as a Measure for Physical Reasoning Intelligence. Nature Machine Intelligence 5, 1 (2023), 83–93.
- Yang et al. (2024) Haoran Yang, Hongyuan Lu, Wai Lam, and Deng Cai. 2024. Exploring Compositional Generalization of Large Language Models. In NAACL Workshop.
- Yao et al. (2023) Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Thomas L Griffiths, Yuan Cao, and Karthik Narasimhan. 2023. Tree of Thoughts: Deliberate Problem Solving with Large Language Models. In NeurIPS.
- Zhang et al. (2019) Chi Zhang, Feng Gao, Baoxiong Jia, Yixin Zhu, and Song-Chun Zhu. 2019. RAVEN: A Dataset for Relational and Analogical Visual Reasoning. In CVPR.
- Zhang et al. (2024b) Jinghan Zhang, Xiting Wang, Weijieying Ren, Lu Jiang, Dongjie Wang, and Kunpeng Liu. 2024b. RATT: A Thought Structure for Coherent and Correct LLM Reasoning. arXiv:2406.02746 (2024).
- Zhang et al. (2024a) Wenqi Zhang, Yongliang Shen, Linjuan Wu, Qiuying Peng, Jun Wang, Yueting Zhuang, and Weiming Lu. 2024a. Self-Contrast: Better Reflection Through Inconsistent Solving Perspectives. In ACL.
- Zhang et al. (2020) Wenhe Zhang, Chi Zhang, Yixin Zhu, and Song-Chun Zhu. 2020. Machine Number Sense: A Dataset of Visual Arithmetic Problems for Abstract and Relational Reasoning. In AAAI.
- Zhou et al. (2023) Denny Zhou, Nathanael Schärli, Le Hou, Jason Wei, Nathan Scales, Xuezhi Wang, Dale Schuurmans, Claire Cui, Olivier Bousquet, Quoc Le, and Ed Chi. 2023. Least-to-Most Prompting Enables Complex Reasoning in Large Language Models. In ICLR.
- Zhou et al. (2022) Hattie Zhou, Azade Nova, Hugo Larochelle, Aaron Courville, Behnam Neyshabur, and Hanie Sedghi. 2022. Teaching Algorithmic Reasoning via In-Context Learning. arXiv:2211.09066 (2022).
Appendix A Supplementary Analysis
A.1. Comparing LLM and Human Problem Difficulty Perception
Following the analysis in Section 3.1.4, we analyzed problems that LLMs (Large Language Models) solve well and those they struggle with. Table 6 presents the accuracy of LLMs across problem difficulty levels classified by humans. The classification was based on the existing categorization relying on perceived difficulty by humans (Borsky, 2021). As a result, we discovered a tendency where problems perceived as difficult by humans align closely with those challenging for LLMs. Difficult problems shared two commonalities: 1) they required lengthy inference processes to solve, and 2) they involved considering multiple simultaneous problems to extract information about changes. An example from Fig. 16 illustrates this point: a task classified as ‘Entry’ only requires a single step of coloring, while a task classified as ‘Hard’ requires three steps: recognizing each object, identifying the priority of each object, and merging each object considering their priority. ‘Easy’ and ‘Medium’ are tasks that require relatively more complex steps than ‘Entry’ and fewer steps than ‘Hard’. Considering these observations, it can be inferred that artificial intelligence possesses simple forms of visual logic that deal with only one of the four priors included in ARC: objectness, goal-directedness, numbers and counting, and basic geometry. However, it cannot handle complex combinations of logic that integrate these priors.
| Entry | Easy | Medium | Hard | |
| Tasks | 2 | 20 | 46 | 14 |
| Trials | 10 | 100 | 230 | 70 |
| CoT | 100.00% | 30.00% | 0.00% | 0.00% |
| LtM | 20.00% | 19.00% | 0.00% | 2.85% |
| ToT | 50.00% | 22.00% | 0.00% | 0.00% |
| Average | 56.67% | 23.67% | 0.00% | 0.95% |
A.2. Comparison of Augmentation Cost-Efficiency across GPT Versions
In a follow-up experiment to our productivity study, we aimed to compare the cost-efficiency of GPT-3.5 and GPT4-32k when augmenting demonstration example tasks. This investigation was crucial to understanding the trade-offs between model performance and associated costs in real-world applications.
Our experimental setup began with the creation of a prompt describing the category, as detailed in Table LABEL:tbl:appendix_productivity_category_prompt. Using this prompt, we developed an Inverse Transformation Prompt (ITP) and proceeded to augment demonstration examples using both GPT-3.5-16k and GPT-4-32k models. Throughout this process, we meticulously logged all prompts given to the LLMs and their corresponding responses.
To analyze the cost implications, we tokenized the logged text using the tiktoken library. We then calculated the cost of generating a single valid demonstration example based on the per-token cost specified by the Azure OpenAI API. This approach allowed us to accurately assess the financial implications of using each model for demonstration example augmentation. Validation of the generated examples was a critical component of our experiment. We employed human reviewers to manually verify the quality and appropriateness of the outputs. These reviewers were tasked with confirming two key aspects:
Whether the results could be legitimately generated from the given rules.
If the generated results were unique, avoid repetition or trivial variations.
This rigorous validation process ensured that our assessment of “valid” examples was thorough and meaningful in the context of practical applications.
| Generated Example | Valid Example | Validity | Cost per Valid Example | |
| GPT-3.5-16k | 346 | 24 | 6.94% | $0.0275 |
| GPT-4-32k | 411 | 40 | 9.73% | $0.3925 |
Analysis of the cost to generate valid demonstration examples, as illustrated in Table 7, reveals that while GPT-4-32k showed approximately 1.5 times higher performance in terms of validity compared to GPT-3.5-16k, its cost was nearly 20 times higher. This suggests that productivity gains may not scale linearly with model capability and cost, especially when generating outputs under complex constraints. Consequently, in scenarios requiring valid outputs under intricate constraints, GPT-3.5 might be preferable to GPT4-32k when considering the trade-off between performance improvement and cost increase. However, the low overall validity rate of less than 10% for both models indicates that current LLMs still have significantly lower productivity compared to humans in such tasks. This finding suggests that merely upgrading to more advanced models is unlikely to fully resolve the productivity gap, highlighting the need for further research and development in enhancing LLM performance for complex, constrained tasks.
Appendix B Experimental Detail
B.1. Logical Coherence
The logical coherence study comprised two main experiments: a comparison on semantic coherence across prompting techniques and an assessment of the inferential coherence of LLMs. For the first experiment, prompting techniques comparison, we randomly selected 100 tasks from the ARC evaluation set. We then applied three different prompting methods - Chain of Thought (CoT), Least to Most (LtM), and Tree of Thoughts (ToT) - to compare their effectiveness in maintaining semantic coherence.
The second experiment assessing the inferential coherence of the LLMs aims to assess whether the same logic can be consistently applied. Therefore, it is necessary to first confirm the tasks where the LLMs have understood the logic. To this end, we experimented using CoT prompting, which showed the best performance in the comparison across prompting techniques experiment, to solve the ARC training set. This experiment was repeated five times. The Inferential Coherence of the LLMs experiment was then conducted on tasks that were correctly solved at least once out of the five repetitions. The detailed task IDs and prompts used in each experiment are provided in B.1.1 and B.1.2, respectively.
B.1.1. Task ID List for Each Experiment
| Experiment | Task ID List |
| Comparison Across Prompting Techniques | ‘3ed85e70’, ‘0f63c0b9’, ‘17cae0c1’, ‘47996f11’, ‘4acc7107’, ‘0692e18c’, ‘477d2879’, ‘1c0d0a4b’, ‘292dd178’, ‘1990f7a8’, ‘22a4bbc2’, ‘4364c1c4’, ‘2f0c5170’, ‘17b80ad2’, ‘03560426’, ‘0c786b71’, ‘3391f8c0’, ‘42a15761’, ‘0bb8deee’, ‘1e97544e’, ‘1c02dbbe’, ‘4b6b68e5’, ‘2a5f8217’, ‘3194b014’, ‘1acc24af’, ‘0c9aba6e’, ‘0e671a1a’, ‘37d3e8b2’, ‘0becf7df’, ‘0607ce86’, ‘3a301edc’, ‘2546ccf6’, ‘009d5c81’, ‘31adaf00’, ‘281123b4’, ‘3d31c5b3’, ‘423a55dc’, ‘1d0a4b61’, ‘1a2e2828’, ‘319f2597’, ‘3979b1a8’, ‘12422b43’, ‘140c817e’, ‘0a2355a6’, ‘19bb5feb’, ‘332efdb3’, ‘27a77e38’, ‘2c0b0aff’, ‘00dbd492’, ‘2c737e3’, ‘2072aba6’, ‘48f8583b’, ‘27f8ce4f’, ‘14754a24’, ‘32e9702f’, ‘195ba7dc’, ‘137f0df0’, ‘184a9768’, ‘29700607’, ‘1c56ad9f’, ‘15663ba9’, ‘4c177718’, ‘136b0064’, ‘0a1d4ef5’, ‘1d398264’, ‘09c534e7’, ‘2685904e’, ‘48131b3c’, ‘31d5ba1a’, ‘2697da3f’, ‘103eff5b’, ‘12997ef3’, ‘1e81d6f9’, ‘25094a63’, ‘08573cc6’, ‘20981f0e’, ‘4852f2fa’, ‘2b01abd0’, ‘2072aba6’, ‘1a6449f1’, ‘34b99a2b’, ‘0b17323b’, ‘15696249’, ‘414297c0’, ‘2753e76c’, ‘12eac192’, ‘0934a4d8’, ‘310f3251’, ‘358ba94e’, ‘21f83797’, ‘4aab4007’, ‘351d6448’, ‘45bbe264’, ‘456873bc’, ‘15113be4’, ‘3490cc26’, ‘3b4c2228’, ‘00576224’, ‘42918530’, ‘45737921’, ‘20818e16’ |
| Inferential Coherence of LLMs | ‘017c7c7b’, ‘025d127b’, ‘08ed6ac7’, ‘0dfd9992’, ‘1bfc4729’, ‘22eb0ac0’, ‘239be575’, ‘23b5c85d’, ‘25d8a9c8’, ‘25ff71a9’, ‘272f95fa’, ‘27a28665’, ‘3618c87e’, ‘3af2c5a8’, ‘3bd67248’, ‘3bdb4ada’, ‘44f52bb0’, ‘48d8fb45’, ‘496994bd’, ‘49d1d64f’, ‘539a4f51’, ‘53b68214’, ‘5582e5ca’, ‘5bd6f4ac’, ‘6150a2bd’, ‘62c24649’, ‘67a3c6ac’, ‘67e8384a’, ‘68b16354’, ‘6d0aefbc’, ‘6f8cd79b’, ‘6fa7a44f’, ‘7447852a’, ‘746b3537’, ‘74dd1130’, ‘7837ac64’, ‘794b24be’, ‘7b7f7511’, ‘7f4411dc’, ‘82819916’, ‘88a62173’, ‘8be77c9e’, ‘8e1813be’, ‘90c28cc7’, ‘9172f3a0’, ‘963e52fc’, ‘97999447’, ‘9dfd6313’, ‘a416b8f3’, ‘a65b410d’, ‘a699fb00’, ‘a85d4709’, ‘a87f7484’, ‘aabf363d’, ‘b1948b0a’, ‘b8cdaf2b’, ‘b94a9452’, ‘ba26e723’, ‘ba97ae07’, ‘bc1d5164’, ‘bd4472b8’, ‘bda2d7a6’, ‘bdad9b1f’, ‘c3f564a4’, ‘c59eb873’, ‘c8f0f002’, ‘c9e6f938’, ‘c9f8e694’, ‘d0f5fe59’, ‘d10ecb37’, ‘d13f3404’, ‘d2abd087’, ‘d4469b4b’, ‘d631b094’, ‘d8c310e9’, ‘d9fac9be’, ‘dc433765’, ‘e26a3af2’, ‘e9afcf9a’, ‘ea786f4a’, ‘ef135b50’, ‘f5b8619d’, ‘f76d97a5’ |
B.1.2. Prompting Setting
The prompts used in comparison across prompting techniques and Inferential Coherence of LLMs are CoT, L2M, and ToT, as shown in Figure 18. These are detailed in Table 9. In the prompts, parts enclosed in curly brackets indicate where the corresponding Type should be inserted. Demo examples and test input refer to the demonstration examples and test input of the test example provided in the task to be solved. For instance, if the type is CoT Prompt, it consists of the CoT one-shot example, the task’s demonstration examples, and test input. Regardless of the prompting method, all prompts are given a one-shot example. CoT solves the task using only the one-shot example, the task’s demonstration examples, and test input. On the other hand, LtM and ToT use a decomposing prompt to obtain instructions for solving the problem through a decomposing stage. For LtM, the step-by-step solving prompt is then used to execute the instructions obtained through decomposing one by one. The previously executed instructions and the resulting changed grid are included in this process. For ToT, the decomposing prompt is used to create multiple instruction candidates, and the ToT decomposing vote prompt is used to have the LLM select the most promising instruction candidate. The selected instructions are then processed through the step-by-step solving prompt to generate multiple candidate results for each instruction. The ToT step-by-step solving vote prompt is then used to select the grid that best reflects the instruction. This process is carried out step-by-step for all instructions.
| Type | Prompt |
| CoT One-Shot Example Data |
If input grids are like that
[[0, 3, 0, 0, 0, 0], [0, 3, 0, 2, 0, 0], [0, 0, 0, 2, 0, 0], [0, 8, 0, 0, 2, 2], [0, 0, 0, 0, 2, 2], [6, 6, 6, 0, 0, 0]], then these grids change to output grids below [[0, 0, 0, 0, 3, 0], [0, 0, 0, 0, 3, 2], [0, 0, 0, 0, 0, 2], [0, 0, 0, 8, 2, 2], [0, 0, 0, 0, 2, 2], [0, 0, 0, 6, 6, 6]]. |
| CoT One-Shot Example |
Do you know ARC problem?
It is similar to below. {CoT One-Shot Example Data} You can understand the pattern of this problem with the input-output pair in Example 1. In the above case, you can infer as follows. In Example 1, all objects move to the right, and then you can move the object to the right side. Like this concept, ‘object’, ‘count’, ‘color’, ‘move’, ‘row’, ‘column’, etc., help you to understand the patterns of the problem and solve it. Below is another pattern to solve. So you can understand the pattern with several examples and apply the problem’s input to get the correct output. |
| CoT Prompt |
{CoT One-Shot Example}
{Demo Examples} If input grids are like that {Test Input} then output grids? |
| Decomposing One-Shot Example-Demo Examples |
Example 1
If input grids are like that [[0, 0, 0, 0, 0, 0], [0, 0, 3, 0, 0, 0], [0, 3, 0, 3, 0, 0], [0, 0, 3, 0, 3, 0], [0, 0, 0, 3, 0, 0], [0, 0, 0, 0, 0, 0]] then these grids change to output grids below [[0, 0, 0, 0, 0, 0], [0, 0, 3, 0, 0, 0], [0, 3, 4, 3, 0, 0], [0, 0, 3, 4, 3, 0], [0, 0, 0, 3, 0, 0], [0, 0, 0, 0, 0, 0]]. Example 2 If input grids are like that [[0, 0, 0, 0, 0, 0, 0, 0, 0, 0], [0, 0, 3, 0, 3, 0, 0, 0, 0, 0], [0, 0, 0, 3, 0, 3, 0, 0, 0, 0], [0, 0, 3, 0, 0, 0, 3, 0, 0, 0], [0, 0, 0, 0, 0, 3, 0, 3, 0, 0], [0, 0, 0, 3, 0, 3, 3, 0, 0, 0], [0, 0, 3, 3, 3, 0, 0, 0, 0, 0], [0, 0, 0, 3, 0, 0, 0, 0, 0, 0], [0, 0, 0, 0, 0, 0, 0, 0, 0, 0], [0, 0, 0, 0, 0, 0, 0, 0, 0, 0]] then these grids change to output grids below [[0, 0, 0, 0, 0, 0, 0, 0, 0, 0], [0, 0, 3, 0, 3, 0, 0, 0, 0, 0], [0, 0, 0, 3, 0, 3, 0, 0, 0, 0], [0, 0, 3, 0, 0, 0, 3, 0, 0, 0], [0, 0, 0, 0, 0, 3, 4, 3, 0, 0], [0, 0, 0, 3, 0, 3, 3, 0, 0, 0], [0, 0, 3, 3, 3, 0, 0, 0, 0, 0], [0, 0, 0, 3, 0, 0, 0, 0, 0, 0], [0, 0, 0, 0, 0, 0, 0, 0, 0, 0], [0, 0, 0, 0, 0, 0, 0, 0, 0, 0]]. |
| Decomposing One-Shot Example-Test Input |
Example-problem
If input grids are like that [[0, 0, 0, 0, 0, 3, 0, 0, 0, 0], [0, 0, 0, 0, 3, 0, 0, 0, 0, 0], [0, 3, 3, 0, 3, 3, 0, 3, 0, 0], [3, 0, 0, 3, 0, 0, 3, 0, 3, 0], [0, 0, 0, 3, 0, 0, 3, 3, 0, 0], [0, 0, 0, 3, 0, 0, 3, 0, 0, 0], [0, 0, 0, 3, 0, 0, 3, 0, 0, 0], [0, 0, 0, 0, 3, 3, 0, 3, 0, 0], [0, 0, 0, 0, 0, 0, 0, 0, 3, 0], [0, 0, 0, 0, 0, 0, 0, 0, 0, 0]] then output grids? |
| Decomposing Prompt |
Do you know the ARC problem?
Each example has the same pattern and the quiz also has the same pattern with examples. So If you understand the pattern of examples, you can solve the quiz. It will help you to analyze the pattern if you decompose the pattern into some steps. I give an example that decomposes patterns into subquestions. {Decomposing One-Shot Example-Demo Examples} {Decomposing One-Shot Example-Test Input} To solve the quiz, I think we should do something like below Q1: We need to identify the places surrounded by 3s in the input grid of example-quiz. Q2: Fill in the 4 in the location we found through Q1. {Demo Examples} {Test Input} I want you to answer in the format below Q1: …. Q2: …. … QN: …. N is the index of the last question. (The answers to the last question should allow you to generate the output grid for the quiz, and you shouldn’t solve the problem yet in this process. You should only create the subquestions for solving the problem.) |
| Step-By-Step Solving Prompt |
{CoT One-Shot Example}
{Demo Examples} If input grids are like that {Test Input} then output grids? {Previous Instructions} {Previous Changed Grid} {Current Instruction} |
| ToT Decomposing Vote Prompt |
First, consider how to solve the problem below.
{Demo Examples} {Test Input} Then, given instruction and several choices, decide which choice is most promising. Analyze each choice in detail, then conclude in the last line ”The best choice is s”, where s is the integer ID of the choice. |
| ToT Step-By-Step Solving Vote Prompt |
First, consider how to solve the problem below.
{Demo Examples} {Test Input} Then, given a question and some answers, you need to choose which answer is the best answer to the question. Analyze each choice in detail and then conclude on the last line, ”The best answer is ‘s’,” where s is the integer ID of the answer. {Previous Instructions} {Previous Changed Grid} {Current Instruction} {Current Changed Grid} |
B.2. Compositionality
In the compositionality study, we conducted two experiments: the LLMs DSL understanding experiment to assess how well LLMs comprehend the provided DSLs, and an experiment to evaluate LLMs’ compositionality ability. The LLMs DSL understanding experiment measures how accurately LLMs can generate the correct DSLs when given the answer for a task. The compositionality ability experiment examines whether LLMs can correctly select and use the necessary DSLs from those provided for problem-solving. Both experiments used the same set of tasks. Detailed information about the task IDs can be found in Table LABEL:tbl:appendix_compositionality_task_id_list, while the specific Prompt details are available in Table LABEL:tbl:appendix_compositionality_dsl_function_prompt and Table LABEL:tbl:appendix_compositionality_prompt_setting.
B.2.1. Task ID List
| Experitment | Task ID List |
| LLMs DSL understanding & Compositionality of LLMs | ‘025d127b’, ‘05f2a901’, ‘08ed6ac7’, ‘0ca9ddb6’, ‘0d3d703e’, ‘11852cab’, ‘150deff5’, ‘1b60fb0c’, ‘1caeab9d’, ‘1e0a9b12’, ‘1f642eb9’, ‘1f876c06’, ‘2204b7a8’, ‘22168020’, ‘22233c11’, ‘228f6490’, ‘22eb0ac0’, ‘253bf280’, ‘25d487eb’, ‘25d8a9c8’, ‘25ff71a9’, ‘29c11459’, ‘2bee17df’, ‘2c608aff’, ‘2dd70a9a’, ‘3345333e’, ‘3618c87e’, ‘36fdfd69’, ‘3906de3d’, ‘3aa6fb7a’, ‘3bd67248’, ‘3c9b0459’, ‘3e980e27’, ‘3eda0437’, ‘40853293’, ‘42a50994’, ‘444801d8’, ‘44d8ac46’, ‘4938f0c2’, ‘496994bd’, ‘50846271’, ‘508bd3b6’, ‘50cb2852’, ‘5168d44c’, ‘54d82841’, ‘5582e5ca’, ‘56dc2b01’, ‘56ff96f3’, ‘5c0a986e’, ‘60b61512’, ‘6150a2bd’, ‘623ea044’, ‘63613498’, ‘67385a82’, ‘673ef223’, ‘67a3c6ac’, ‘67a423a3’, ‘6855a6e4’, ‘68b16354’, ‘694f12f3’, ‘6c434453’, ‘6cf79266’, ‘6d58a25d’, ‘6d75e8bb’, ‘6e02f1e3’, ‘6e19193c’, ‘6e82a1ae’, ‘74dd1130’, ‘760b3cac’, ‘776ffc46’, ‘794b24be’, ‘7ddcd7ec’, ‘7e0986d6’, ‘7f4411dc’, ‘810b9b61’, ‘855e0971’, ‘88a10436’, ‘890034e9’, ‘8d510a79’, ‘90f3ed37’, ‘928ad970’, ‘93b581b8’, ‘941d9a10’, ‘952a094c’, ‘9565186b’, ‘99fa7670’, ‘9dfd6313’, ‘a2fd1cf0’, ‘a3df8b1e’, ‘a48eeaf7’, ‘a5313dff’, ‘a61f2674’, ‘a65b410d’, ‘a699fb00’, ‘a78176bb’, ‘a79310a0’, ‘a85d4709’, ‘a9f96cdd’, ‘aabf363d’, ‘ae3edfdc’, ‘aedd82e4’, ‘af902bf9’, ‘b1948b0a’, ‘b230c067’, ‘b2862040’, ‘b548a754’, ‘b7249182’, ‘b8cdaf2b’, ‘ba97ae07’, ‘bb43febb’, ‘bda2d7a6’, ‘bdad9b1f’, ‘c0f76784’, ‘c8f0f002’, ‘cbded52d’, ‘ce22a75a’, ‘ce9e57f2’, ‘d037b0a7’, ‘d07ae81c’, ‘d23f8c26’, ‘d2abd087’, ‘d406998b’, ‘d43fd935’, ‘d4a91cb9’, ‘d4f3cd78’, ‘d511f180’, ‘d5d6de2d’, ‘d6ad076f’, ‘d89b689b’, ‘d8c310e9’, ‘d90796e8’, ‘d9f24cd1’, ‘dc433765’, ‘ddf7fa4f’, ‘ded97339’, ‘e40b9e2f’, ‘e48d4e1a’, ‘e5062a87’, ‘e509e548’, ‘e73095fd’, ‘e8dc4411’, ‘e9614598’, ‘e9afcf9a’, ‘ea32f347’, ‘ea786f4a’, ‘ec883f72’, ‘ed36ccf7’, ‘ef135b50’, ‘f25ffba3’, ‘f76d97a5’, ‘f8a8fe49’, ‘fcc82909’, ‘8f2ea7aa’, ‘5521c0d9’, ‘32597951’, ‘98cf29f8’, ‘0e206a2e’, ‘a1570a43’ |
B.2.2. Types of DSLs used
Each DSL was implemented as a Python function. As shown in Table 12, there are three types of DSLs using three parameter types. Color Change DSLs accept parameters such as Coordinate and Object. Coordinate-based Color Change DSLs include pixel color, X line, horizontal line, vertical line, and diagonal line. For Object parameters, only the obj color DSL exists. Transformation DSLs use Object and Grid parameters. Object-based transformations include rotate left obj, rotate right obj, horizontal flip obj, vertical flip obj, and movement operations (move left, right, up, down). Grid-based transformations include rotate left state, rotate right state, horizontal flip, and vertical flip. Lastly, the Complete DSL exists independently of parameters, indicating task completion before reaching 10 steps. For tasks solved in exactly 10 steps, the Complete DSL is unnecessary.
| Coordinate | Object | Grid | |||||||||
| Color Change |
obj color | X | ||||||||
| Transformation | X |
| Complete | Complete | ||||||||||
B.2.3. Prompt Contents for LLMs with DSL Codes and Comments
In the two experiments measuring compositionality and LLM’s DSL understanding, we identified a set of 10 tasks that collectively required the use of all 15 DSLs at least once. This set was used to determine the optimal prompt for explaining DSLs to LLMs. We conducted experiments with four prompt variants: no DSL information, DSL code only, DSL comments only, and both DSL code and comments. The LLMs DSL understanding experiment was performed for these 10 tasks across all four prompt compositions. Results indicated that providing both code and comments yielded optimal performance. Consequently, for both the LLMs DSL understanding and compositionality of LLMs experiments, we employed prompts containing both DSL code and comments. Table LABEL:tbl:appendix_compositionality_dsl_function_prompt illustrates the prompt content where both code and comments were provided to the LLM.
| Type | Prompt |
| Rotate Left State |
# rotate_left_state function is a counterclockwise rotation about the given state.
# This function rotates a square grid (NxN) counterclockwise by 90 degrees. # Parameters: # - state: A 2D list representing the current grid state. # Returns: # - A new 2D list representing the grid after the counterclockwise rotation. def rotate_left_state(state): N = len(state) rotated_state = copy.deepcopy(state) if N == len(state[0]): temp_state = copy.deepcopy(state) for x in range(N): for y in range(N): rotated_state[N-1-y][x] = state[x][y] return rotated_state |
| Rotate Right State |
# rotate_right_state function is a clockwise rotation about the given state.
# This function rotates a square grid (NxN) clockwise by 90 degrees. # Parameters: # - state: A 2D list representing the current grid state. # Returns: # - A new 2D list representing the grid after the clockwise rotation. def rotate_right_state(state): N = len(state) rotated_state = copy.deepcopy(state) if N == len(state[0]): for x in range(N): for y in range(N): rotated_state[y][N-1-x] = state[x][y] return rotated_state |
| Vertical Flip |
# vertical_flip function is a flip by x-axis about the given state.
# This function flips the grid vertically, swapping the top and bottom rows. # Parameters: # - state: A 2D list representing the current grid state. # Returns: # - A new 2D list representing the grid after the vertical flip. def vertical_flip(state): temp_state = copy.deepcopy(state) N = len(state) M = len(state[0]) for i in range(N): for j in range(M): temp_state[N-1-i][j] = state[i][j] return temp_state |
| Horizontal Flip |
# horizontal_flip function is a flip by y-axis about the given state.
# This function flips the grid horizontally, swapping the left and right columns. # Parameters: # - state: A 2D list representing the current grid state. # Returns: # - A new 2D list representing the grid after the horizontal flip. def horizontal_flip(state): N = len(state) M = len(state[0]) flipped_state = copy.deepcopy(state) for i in range(N): for j in range(M // 2): flipped_state[i][j], flipped_state[i][M-1-j] = state[i][M-1-j], state[i][j] return flipped_state |
| Move Right |
# move_right function moves all pixels in the selected object to the right by one column.
# Parameters: # - state: A 2D list representing the current grid state. # - object: A list of lists where each inner list contains the coordinates [x, y] of a pixel to move. # Returns: # - A new 2D list representing the grid after the object is moved to the right. def move_right(state, object): move_state = copy.deepcopy(state) new_obj = [] for x, y in object: move_state[x][y] = 0 for x, y in object: new_x, new_y = x, y + 1 if 0 ¡= new_x ¡ len(state) and 0 ¡= new_y ¡ len(state[0]): move_state[new_x][new_y] = state[x][y] new_obj.append([new_x, new_y]) return move_state |
| Move Left |
# move_left function moves all pixels in the selected object to the left by one column.
# Parameters: # - state: A 2D list representing the current grid state. # - object: A list of lists where each inner list contains the coordinates [x, y] of a pixel to move. # Returns: # - A new 2D list representing the grid after the object is moved to the left. def move_left(state, object): move_state = copy.deepcopy(state) new_obj = [] for x, y in object: move_state[x][y] = 0 for x, y in object: new_x, new_y = x, y - 1 if 0 ¡= new_x ¡ len(state) and 0 ¡= new_y ¡ len(state[0]): move_state[new_x][new_y] = state[x][y] new_obj.append([new_x, new_y]) return move_state |
| Move Up |
# move_up function moves all pixels in the selected object up by one row.
# Parameters: # - state: A 2D list representing the current grid state. # - object: A list of lists where each inner list contains the coordinates [x, y] of a pixel to move. # Returns: # - A new 2D list representing the grid after the object is moved up. def move_up(state, object): move_state = copy.deepcopy(state) new_obj = [] for x, y in object: move_state[x][y] = 0 for x, y in object: new_x, new_y = x-1, y if 0 ¡= new_x ¡ len(state) and 0 ¡= new_y ¡ len(state[0]): move_state[new_x][new_y] = state[x][y] new_obj.append([new_x, new_y]) return move_state |
| Move Down |
# move_down function moves all pixels in the selected object down by one row.
# Parameters: # - state: A 2D list representing the current grid state. # - object: A list of lists where each inner list contains the coordinates [x, y] of a pixel to move. # Returns: # - A new 2D list representing the grid after the object is moved down. def move_down(state, object): move_state = copy.deepcopy(state) new_obj = [] for x, y in object: move_state[x][y] = 0 for x, y in object: new_x, new_y = x+1, y if 0 ¡= new_x ¡ len(state) and 0 ¡= new_y ¡ len(state[0]): move_state[new_x][new_y] = state[x][y] new_obj.append([new_x, new_y]) return move_state |
| Rotate Right Object |
# rotate_right_obj function makes a clockwise rotation about the given object.
# This function rotates the selected object within the grid 90 degrees clockwise around its center. # Parameters: # - state: A 2D list representing the current grid state. # - object: A list of lists where each inner list contains the coordinates [x, y] of a pixel in the object. # Returns: # - A new 2D list representing the grid after the object is rotated clockwise. def rotate_right_obj(state, object): rotate_state = copy.deepcopy(state) new_obj = [] max_x = max(x for x,_ in object) min_x = min(x for x,_ in object) max_y = max(y for_, y in object) min_y = min(y for_, y in object) fixed_x = (max_x + min_x) // 2 fixed_y = (max_y + min_y) // 2 for x, y in object: rotate_state[x][y] = 0 for x, y in object: moved_x = y - fixed_y + fixed_x moved_y = -x + fixed_x + fixed_y if 0 ¡= moved_x ¡ len(state) and 0 ¡= moved_y ¡ len(state[0]): rotate_state[moved_x][moved_y] = state[x][y] new_obj.append([moved_x, moved_y]) return rotate_state |
| Rotate Left Object |
# rotate_left_obj function makes a counterclockwise rotation about the given object.
# This function rotates the selected object within the grid 90 degrees counterclockwise around its center. # Parameters: # - state: A 2D list representing the current grid state. # - object: A list of lists where each inner list contains the coordinates [x, y] of a pixel in the object. # Returns: # - A new 2D list representing the grid after the object is rotated counterclockwise. def rotate_left_obj(state, object): rotate_state = copy.deepcopy(state) new_obj = [] max_x = max(x for x,_ in object) min_x = min(x for x,_ in object) max_y = max(y for_, y in object) min_y = min(y for_, y in object) fixed_x = (max_x + min_x) // 2 fixed_y = (max_y + min_y) // 2 for x, y in object: rotate_state[x][y] = 0 for x, y in object: moved_x = -y + fixed_y + fixed_x moved_y = x - fixed_x + fixed_y if 0 ¡= moved_x ¡ len(state) and 0 ¡= moved_y ¡ len(state[0]): rotate_state[moved_x][moved_y] = state[x][y] new_obj.append([moved_x, moved_y]) return rotate_state |
| Vertical Flip Object |
# vertical_flip_obj function makes a vertical flip of the selected object.
# This function flips the selected object within the grid vertically. # Parameters: # - state: A 2D list representing the current grid state. # - object: A list of lists where each inner list contains the coordinates [x, y] of a pixel in the object. # Returns: # - A new 2D list representing the grid after the object is flipped vertically. def vertical_flip_obj(state, object): flip_state = copy.deepcopy(state) new_obj = [] max_x = max(x for x,_ in object) min_x = min(x for x,_ in object) for x, y in object: flip_state[x][y] = 0 for x, y in object: flip_state[max_x + min_x - x][y] = state[x][y] new_obj.append([max_x + min_x - x, y]) return flip_state |
| Horizontal Flip Object |
# horizontal_flip_obj function makes a horizontal flip of the selected object.
# This function flips the selected object within the grid horizontally. # Parameters: # - state: A 2D list representing the current grid state. # - object: A list of lists where each inner list contains the coordinates [x, y] of a pixel in the object. # Returns: # - A new 2D list representing the grid after the object is flipped horizontally. def horizontal_flip_obj(state, object): flip_state = copy.deepcopy(state) new_obj = [] max_y = max(y for_, y in object) min_y = min(y for_, y in object) for x, y in object: flip_state[x][y] = 0 for x, y in object: flip_state[x][max_y + min_y - y] = state[x][y] new_obj.append([x, max_y + min_y - y]) return flip_state |
| X Line |
# X_line function makes a diagonal X-line in all directions from a given pixel until they reach the end of the grid.
# Parameters: # - state: A 2D list representing the current grid state. # - r: The row index of the starting pixel. # - c: The column index of the starting pixel. # - color: The color to be used for the X-line. # Returns: # - A new 2D list representing the grid after the X-line is drawn. def X_line(state, r, c, color): X_state = copy.deepcopy(state) x_move = [-1, 1] y_move = [-1, 1] for i in x_move: for j in y_move: moved_x, moved_y = r + i, c + j while 0 ¡= moved_x ¡ len(state) and 0 ¡= moved_y ¡ len(state[0]): X_state[moved_x][moved_y] = color moved_x += i moved_y += j return X_state |
| Horizontal Line |
# horizontal_line function draws a horizontal line between two pixels if they are on the same row.
# Parameters: # - state: A 2D list representing the current grid state. # - r1: The row index of the first pixel. # - c1: The column index of the first pixel. # - r2: The row index of the second pixel. # - c2: The column index of the second pixel. # - color: The color to be used for the line. # Returns: # - A new 2D list representing the grid after the horizontal line is drawn. def horizontal_line(state, r1, c1, r2, c2, color): line_state = copy.deepcopy(state) if r1 == r2: if c1 ¡ c2: if c2 ¡ len(state[0]): for i in range(c1+1, c2): line_state[r1][i] = color else: if c1 ¡ len(state[0]): for i in range(c2+1, c1): line_state[r1][i] = color return line_state |
| Vertical Line |
# vertical_line function draws a vertical line between two pixels if they are in the same column.
# Parameters: # - state: A 2D list representing the current grid state. # - r1: The row index of the first pixel. # - c1: The column index of the first pixel. # - r2: The row index of the second pixel. # - c2: The column index of the second pixel. # - color: The color to be used for the line. # Returns: # - A new 2D list representing the grid after the vertical line is drawn. def vertical_line(state, r1, c1, r2, c2, color): line_state = copy.deepcopy(state) if c1 == c2: if r1 ¡ r2: if r2 ¡ len(state): for i in range(r1+1, r2): line_state[i][c1] = color else: if r1 ¡ len(state): for i in range(r2+1, r1): line_state[i][c1] = color return line_state |
| Diagonal Line |
# diagonal_line function draws a diagonal line between two pixels if they form a 45-degree angle.
# Parameters: # - state: A 2D list representing the current grid state. # - r1: The row index of the first pixel. # - c1: The column index of the first pixel. # - r2: The row index of the second pixel. # - c2: The column index of the second pixel. # - color: The color to be used for the line. # Returns: # - A new 2D list representing the grid after the diagonal line is drawn. def diagonal_line(state, r1, c1, r2, c2, color): line_state = copy.deepcopy(state) if abs(r1 - r2) == abs(c1 - c2): dr = 1 if r2 ¿ r1 else -1 dc = 1 if c2 ¿ c1 else -1 r, c = r1 + dr, c1 + dc while r != r2 and c != c2: line_state[r][c] = color r += dr c += dc return line_state |
| Object Color |
# obj_color function changes the color of the selected object.
# Parameters: # - state: A 2D list representing the current grid state. # - object: A list of lists where each inner list contains the coordinates [x, y] of a pixel in the object. # - color: The new color to be applied to the object. # Returns: # - A new 2D list representing the grid after the object’s color is changed. def obj_color(state, object, color): color_state = copy.deepcopy(state) for x, y in object: color_state[x][y] = color return color_state |
| Pixel Color |
# pixel_color function changes the color of the selected pixel.
# Parameters: # - state: A 2D list representing the current grid state. # - r: The row index of the pixel to change. # - c: The column index of the pixel to change. # - color: The new color to be applied to the pixel. # Returns: # - A new 2D list representing the grid after the pixel’s color is changed. def pixel_color(state, r, c, color): temp_state = copy.deepcopy(state) temp_state[r][c] = color return temp_state |
| Complete |
# complete function returns the current state as the final answer of the quiz.
# Parameters: # - state: A 2D list representing the current grid state. # Returns: # - The same 2D list representing the final grid state. def complete(state): return state |
B.2.4. Prompt of Compositionality Experiment
Both the LLMs DSL understanding and compositionality of LLMs experiments utilized the prompt structure outlined in Table LABEL:tbl:appendix_compositionality_prompt_setting. The Introduction ARC Prompt provides a comprehensive overview of ARC, while the DSL Usage Example Prompt illustrates DSL application. The DSL prompt, comprising the prompts of DSL function codes and comments from Table LABEL:tbl:appendix_compositionality_dsl_function_prompt and the DSL usage example prompt, offers a comprehensive DSL explanation. The task prompt includes demonstration examples, test input, object information (coordinates of objects obtained through PnP in dictionary format), and output format guidelines. In the case of the LLMs DSL understanding prompt, unlike the task prompt, the DSLs path for the task is provided. The CoT prompt, included the introduction ARC prompt and DSL prompt. In the case of the LLMs DSL understanding experiment, the LLMs DSL understanding prompt was used, while in the case of the compositionality of LLMs experiment, the task prompt was used. In the compositionality experiments, the CoT prompt was utilized.
| Type | Prompt |
| Introduction ARC Prompt |
Do you know ARC problem?
ARC is a quiz, and if we can solve this problem, we can understand and utilize several concepts such as ”object”, ”count”, ”color”, ”move”, ”row”, ”column”, etc. ARC problems give you some examples to understand these patterns. You can understand the pattern below with several examples and then apply the quiz’s input to get the correct output. |
| DSL Usage Example Prompt |
In this example grid,
[[0, 0, 0, 0, 0, 0, 0, 0, 0, 0], [0, 0, 0, 0, 0, 0, 0, 5, 5, 0], [0, 5, 5, 0, 0, 0, 0, 5, 5, 0], [0, 0, 5, 5, 0, 0, 0, 0, 0, 0], [0, 0, 0, 0, 0, 0, 0, 0, 0, 0], [0, 0, 0, 0, 0, 0, 0, 0, 0, 5], [0, 0, 0, 0, 0, 5, 5, 0, 0, 5], [0, 5, 0, 0, 0, 0, 0, 0, 0, 5], [0, 5, 0, 0, 5, 0, 0, 0, 0, 0], [0, 0, 0, 5, 5, 0, 0, 0, 0, 0]] there are 6 objects Object1: [[1, 7], [1, 8], [2, 7], [2, 8]] Object2: [[2, 1], [2, 2], [3, 2], [3, 3]] Object3: [[5, 9], [6, 9], [7, 9]] Object4: [[6, 5], [6, 6]] Object5: [[7, 1], [8, 1]] Object6: [[8, 4], [9, 3], [9, 4]] If you apply ”rotate_right_obj(state, object2)”, the result becomes [[0, 0, 0, 0, 0, 0, 0, 0, 0, 0], [0, 0, 5, 0, 0, 0, 0, 5, 5, 0], [0, 5, 5, 0, 0, 0, 0, 5, 5, 0], [0, 5, 0, 0, 0, 0, 0, 0, 0, 0], [0, 0, 0, 0, 0, 0, 0, 0, 0, 0], [0, 0, 0, 0, 0, 0, 0, 0, 0, 5], [0, 0, 0, 0, 0, 5, 5, 0, 0, 5], [0, 5, 0, 0, 0, 0, 0, 0, 0, 5], [0, 5, 0, 0, 5, 0, 0, 0, 0, 0], [0, 0, 0, 5, 5, 0, 0, 0, 0, 0]] |
| DSL Prompt |
{Prompts of DSL Function Codes and Comments}
Arguments for the DSLs are mainly ”state” and ”object”, but some require ”color”, ”row”, ”column”, etc. ”state” is the current state of the grid, which is the entire grid. ”object” is the list of coordinates of the object; there may be multiple objects in the grid, but no DSL requires multiple objects. ”color” is the color of the pixel in the grid, which is a number between 0 and 9. ”row” and ”column” are the coordinate numbers of a pixel in the grid. You can choose from here and apply the DSL to solve the problem. You must input the appropriate arguments to the DSL, or it will not work. {DSL Usage Example Prompt} Please choose the DSL from the list above and provide the proper arguments to solve the problem. |
| Task Prompt |
{Demo Examples} & {Test Input} & {Object Info}
You must solve the given quiz within 10 steps! Select one DSL with proper arguments in each step. If you think the current state is correct, you can select the ”complete” DSL. I want you to answer in the format below. The output should be in the following JSON format: { ‘step’: ”(current_step)”, ‘dsl’: ”(dsl with the arguments for the DSL)”, ‘description’: ”(why you chose this DSL?)” } |
| LLMs DSL Understanding Prompt |
{Demo Examples} & {Test Input} & {Object Info} & {DSLs Path}
Generate the grid step by step by applying the DSLs in the ‘dsls_path’ to the input grid. The output should be in the following JSON format: { ‘step’: ‘Step number’ ‘dsl’: ‘DSL used from the dsls_path to make the final output grid’ ‘grid’: ‘Final output grid’ ‘objects’: ‘Objects in the final output grid’ } |
| CoT Prompt |
{Introduction ARC Prompt}
{DSL Prompt} {LLMs DSL Understanding Prompt} # It is used when conducting LLMs DSL understanding experiment. {Task Prompt} # It is used when conducting compositionality of LLMs experiment. |
B.3. Productivity
In the productivity experiment, we aimed to augment the task’s demonstration example pairs using the Inverse Transformation Prompt (ITP). The ITP consists of a category prompt, which contains a description of the category, task examples, and the target output to be augmented. The detailed structure of the category prompt is provided in Table LABEL:tbl:appendix_productivity_category_prompt, and the structure of the ITP is outlined in Table LABEL:tbl:appendix_productivity_prompt_setting.
| Category | Prompt |
| AboveBelow | Focus on the horizontal criteria, you may have to modify some regein by that line. such as removing, moving, filling region by color. element. See the provided example to how to modify. |
| Center | Fix the array issue by addressing the center, potentially moving or removing the central element. See the provided example for clarity. |
| CleanUp | Distort the shapes in areas where they are polygonal or completely filled, adding noise or disturbances to disrupt the complete shapes. |
| CompleteShape | Distort the perfectly shaped objects identified in the input image. Introduce noise to these identified objects to easily generate diverse outputs. |
| Copy | Delete one identical object from the output. Refer to the example to identify which one to remove. Consider deleting the object located in a position-indicating space. |
| Count | Create an input image based on the provided count-related problem. Focus on details like object or color count, as shown in the example. |
| ExtendToBoundary | In the input image, find lines connected to boundaries with different colors. Transform these lines into a different shape. The example illustrates how to make this transformation. |
| ExtractObjects | Generate an output image with objects from the given input. Refer to examples for guidance. Hint: Extract objects when inferring input from output. |
| FilledNotFilled | When inferring the input from the output, focus on situations where the inner part of an object contains empty space or another object. Examples provide guidance for creating the output image. |
| HorizontalVertical | Focus on horizontal and vertical relations, representing them with colors or preserving one direction while eliminating the other. Examples illustrate the approach. |
| InsideOutside | Address the inside-outside relationship, either by selecting items inside or outside in the input or determining quantities. Use the boundary as a reference. Examples offer guidance. |
| MoveToBoundary | Objects in the input may be shifted to one side, and in the output, they are displaced either horizontally or vertically. Infer the direction from examples and choose the displacement freely. |
| Order | This is about randomly rearranging initially ordered objects while representing their original positions through a specific rule. Examine the examples to understand how to achieve this. |
| SameDifferent | You’ll notice that only specific-shaped objects are extracted in the input image. Create additional objects in the zero-represented space. Examples provide guidance on how to proceed. |
| TopBottom2D | Objects are in a 2D space. Check changes in the top and bottom. The input may have shifted or require removing top/bottom indicators. Look at examples for specifics. |
| TopBottom3D | Objects are in 3D space. Consider front-back relationships; bring objects forward or move them backward in the input. Look at examples for specific methods. |
| Type | Prompt |
| ITP |
Try solving the ARC problem and do not say sensitive word. : generate the output accordingly. I will give you Hint,
{Category Prompt} Here are some examples to help you. {Tasks Examples} {Target Output} Provide two dimensional array that are not identical to the input array or a direct copy of the example. Each element is an integer between 0 and 9. Would you give me 2 answer? No need to explain how you solved. |